### v2

# NASP: Network Slice as a Service Platform for 5G Networks

Abstract
With the rapid global adoption of fifth-generation (5G) mobile telecommunications, the demand for highly flexible private networks has surged. A key beyond‑5G feature is network slicing, where the 3rd Generation Partnership Project (3GPP) defines three main use cases: massive Machine-Type Communications (mMTC), enhanced Mobile Broadband (eMBB), and Ultra-Reliable Low-Latency Communications (URLLC), along with their associated management functions. Similarly, the European Telecommunications Standards Institute (ETSI) provides the Zero-Touch Network and Service Management (ZSM) standard, enabling operation without human intervention. However, current technical documents lack definitions for end-to-end (E2E) management and integration across domains and subnet instances. We present a Network Slice as a Service Platform (NASP) that is agnostic to 3GPP and non-3GPP networks, addressing this gap. The NASP architecture comprises (i) onboarding requests for new slices at the business level, translating them into definitions of physical instances and interfaces among domains, (ii) a hierarchical orchestrator coordinating management functions, and (iii) communication interfaces with network controllers. Our NASP prototype is developed based on technical documents from 3GPP and ETSI, analyzing design overlaps and gaps across different perspectives. Results demonstrate the platform’s adaptability in handling diverse requests via the Communication Service Management Function. Evaluation indicates that Core configuration accounts for 68% of the time required to create a Network Slice Instance. Tests reveal a 93% reduction in data session establishment time when comparing URLLC and Shared scenarios. Finally, we present cost variations for operating the platform with the orchestration of five and ten slices, showing a 112% variation between Edge and Central deployments.
keywords:
5G , Network slicing , Network Slice as a Service Platform (NASP) , End-to-end (E2E) management , Hierarchical orchestrator , Edge computing
journal: Journal of Network and Computer Applications
[…]
[inst1]organization=Universidade do Vale do Rio dos Sinos - UNISINOS,addressline=Cristo Rei, city=São Leopoldo, postcode=93.022-750, state=Rio Grande do Sul, country=Brazil
[…]
[inst2]organization=Instituto Nacional de Telecomunicações (INATEL),addressline=Av. João de Camargo, 510 - Centro, city=Santa Rita do Sapucaí, state=MG, postcode=37536-001, country=Brazil
1 Introduction
The rapid, sustained increase in mobile data traffic, combined with the emerging diversity of service requirements from Massive Machine-Type Communication (mMTC) terminals and vertical enterprises, challenges the traditional mobile‐broadband paradigm (Navarro-Ortiz2020). Legacy networks, initially designed with only a limited set of parameters (e.g., priority class and Quality of Service (QoS) for broadband services), now face the task of accommodating heterogeneous and, in many cases, conflicting operational needs (Vila2020). For instance, a railway network demands seamless, high-speed mobility management along predefined routes, whereas an electricity metering system requires support for static, low-volume transmissions. Recent advances in Network Functions Virtualization (NFV) and Software-Defined Networking (SDN) have paved the way for innovations such as Network Slicing (NS), which enables the logical separation of network functions and resources tailored to specific technical and commercial requirements (Wijethilaka2021; Ksentini2020; Mahdi2023).
NS has conceptual roots in earlier technologies, such as the 802.1Q virtual local‑area network standard (IEEE8021Q-2018) and Internet Engineering Task Force (IETF) RFC 4026 Virtual Private Networks (rfc4026), which provided isolated broadcast and session domains. However, deploying NS in cellular environments introduces additional challenges, including mobility management, control‑plane authentication, and session and charging management in the user plane (mahyoub2024security). Studies led by the Third Generation Partnership Project (3GPP) and initiatives by major industry players indicate an evolution toward End-to-End (E2E) network slices that integrate both 3GPP and non‑3GPP domains, enhancing connectivity for the Internet of Things (IoT) and enabling diverse service customization (Nguyen2016; bertenyi2018ng; Linnartz2022-lb). Despite these advancements, the complexity stemming from distributed interfaces across Radio Access Network (RAN), Transport Network (TN), and Core Network (CN) necessitates novel orchestration and management solutions (Wyszkowski2024; 7962822).
The evolution toward heterogeneous mobile networks, driven by massive IoT deployments, emerging business-oriented services, and stringent performance requirements, motivates the need for a comprehensive, automated approach to NS. Existing studies have addressed resource allocation using machine‑learning techniques (Jiang; Li2018DeepRL) and have developed frameworks for NS control and assurance (Abbas2020IBNSlicingIN; Theodorou; Bega). However, fully integrated E2E orchestration that translates Service Level Agreement (SLA)/ Service Level Specification (SLS) requirements into practical NS designs remains an open challenge, as clear guidelines that bridge the gap between business-level intents and concrete slice configurations spanning the RAN, TN, and CN domains are still lacking in the literature.
The central research question addressed in this work is how standardized components can be orchestrated and integrated to provide Network Slice as a Service (NSaaS). We argue that effective orchestration and integration must be achieved through well‑defined interfaces among diverse network components and a clear hierarchy of responsibilities within the orchestrator. In this context, we propose Network Slice as a Service Platform (NASP), which (i) translates business‑rule templates into subnet‑instance descriptors, (ii) applies hierarchical slice‑deployment strategies, and (iii) integrates closed‑loop management functions by leveraging Continuous Integration/Continuous Delivery (CI/CD) and Infrastructure as Code (IaC) practices. NASP manages E2E slices across both 3GPP and non‑3GPP networks seamlessly.
The main contributions of this work are threefold. First, we introduce NASP, an automated NSaaS platform that translates business‐rule templates into multi‐domain slice descriptors and orchestrates E2E deployments across both 3GPP and non‐3GPP networks. Second, we propose a hierarchical orchestration model, including Communication Service Management Function (CSMF), Network Slice Management Function (NSMF), and domain‐specific Network Slice Subnet Management Functions, which integrate closed‐loop quality assurance via CI/CD and IaC practices, enabling dynamic SLA compliance through AI‐driven telemetry analytics. Third, we develop and release a fully functional prototype built on Kubernetes, Open Network Operating System (ONOS), my5G-RANTester (Silveira2022), and Free 5G Core (Free5gc), demonstrating seamless integration, extensibility, and native support for physical and virtual resources in heterogeneous deployment environments.
Our comprehensive evaluation, covering mMTC, Ultra-Reliable Low Latency Communications (URLLC), shared, and non-3GPP use cases, yields several impactful findings. We observe that 66% of the E2E slice instantiation time is spent on Core domain configuration, pinpointing critical optimization targets. URLLC slices achieve up to a 93% reduction in data-session establishment time compared to shared slices. A prototype closed-loop quality assurance module demonstrates timely detection and mitigation of performance anomalies in our testbed. Finally, a cost-efficiency analysis reveals a 112% variation in monthly operational expenses between edge and centralized deployments, highlighting the economic differences between the two deployment options.
The remainder of this article is organized as follows. Section 2 provides essential background on NFV, SDN, and NS technologies. Section 3 reviews related work and highlights existing gaps in E2E NS orchestration. Section 4 details the proposed NASP architecture and its key design choices. Section 5 describes the implementation and prototype development of the platform. Section 6 outlines the methodology used to assess platform performance, while Section 7 presents the experimental results and analysis. Finally, Section 8 concludes the article and discusses directions for future research.
2 Background and Evolution of 5G Networks
The evolution of mobile telecommunications toward Fifth Generation (5G) introduced network programmability and NSaaS to better meet emerging service and market needs (3gpp.21.915; afolabi2018network). According to 3gpp.23.501, the 5G system and its Service-Based Architecture (SBA) were first specified in Releases 15–16, with continued refinements in later 3GPP releases (TS 23.501 v19.2.0, Rel-19). The architecture retains core elements, User Equipment (UE), Next Generation Radio Access Network (NG-RAN), and 5G Core (5GC), and separates control and user planes via Access and Mobility Management Function (AMF) and User Plane Function (UPF). Figure 1 highlights supporting functions such as Policy Control Function (PCF) (policy), Authentication Server Function (AUSF) (authentication), Network Slice Selection Function (NSSF) (slice selection), Network Data Analytics Function (NWDAF) (analytics), and Network Exposure Function (NEF) (exposure).
Figure 1: 5G Core architecture.
NS is a central 5G innovation, allowing multiple logical networks to share a common infrastructure while meeting distinct SLAs. Each slice orchestrates resources across the RAN, TN, and CN and is characterized according to 3GPP and Global System for Mobile Communications Association (GSMA) guidance (3gpp.23.501; gsma-116) (GSMA NG.116 v10.0). The Network Slice Type (NEST) concept groups key attributes, throughput, latency, and reliability, into a service profile. For example, an Enhanced Mobile Broadband (eMBB) slice may target 99.999% availability, multimedia telephony, specific QoS, and strong session continuity. Moreover, isolation is essential since congestion or scaling in one slice must not degrade others, and separation must hold for performance, management, and security/privacy (3gpp.28.552). In this case, the system can be viewed in three domains, as can be seen in Figure 1: the RAN (gNB with Radio Unit (RU)/ Distributed Unit (DU)/ Centralized Unit (CU)), the TN (access–core transport), and the CN (interfacing with external Data Networks via the UPF).
Releases 17–18 (3gpp.21.917; 3gpp.21.918) extend slicing toward NSaaS, enabling operators to offer tenants customized, on-demand slices. NSaaS exposes management capabilities via standardized Application Programming Interfaces from 3GPP and TM Forum frameworks (gsma-116). Slice management relies on CSMF, NSMF, and NSSMF to translate tenant demands into an E2E deployment. Figure 2 outlines this orchestration, which includes an NS communication interface and a management domain comprising the NSMF, CSMF, and domain-specific entities (RAN-NSSMF, TN-NSSMF, 5GC-NSSMF). These coordinate instantiation, monitoring, and resource allocation across the NG-RAN, TN, and 5GC. Multiple slices (1–N) can coexist with dedicated resources and connect to external Data Networks through defined channels.
Figure 2: NS orchestration architecture.
NS lifecycle management, from preparation (blueprints, reservations) to operation (activation, monitoring, capacity planning) and decommissioning, benefits from modularity and isolation (5gcorebook-ns). The resulting operational complexity motivates Zero-touch Network and Service Management (ZSM), which applies automation, Machine Learning (ML), and Artificial Intelligence (AI) for self-configuration, self-monitoring, self-healing, and self-optimization (etsi-zsm). Therefore, beyond-5G systems build on these principles to deliver agile, customized services that meet diverse SLAs. Building on these advances, 3gpp.21.917; 3gpp.21.918 describe the latest 3GPP specifications (TS 28.541, TS 28.531, TS 28.622, TS 28.623; Releases 18–19), and together with etsi-zsm they outline a precise Network Resource Model (NRM) for standardized management of 5GC, NG-RAN, and NS. Function Information Object Classes, such as AMFFunction, SMFFunction, UPFFunction, GNBDUFunction, GNBCUCPFunction, encapsulates logical behaviors with well-defined attributes, states, and relationships. The subnetwork IOC provides a hierarchical container to organize ManagedElements into administrative domains. For slicing, NetworkSlice and NetworkSliceSubnet IOCs are augmented with ServiceProfile and SliceProfile data types that represent performance, coverage, capacity, and isolation requirements, aligning with GSMA Generic Network Slice Template (GST)/NG.116 service descriptors to maintain globally consistent SLA representations. These definitions are bound to solution sets in Extensible Markup Language (XML), JavaScript Object Notation (JSON), and YANG, supporting interoperability across multi-vendor environments and aligning with European Telecommunications Standards Institute (ETSI) ZSM principles of zero-touch automation.
Releases 18–19 (3gpp.21.918; 3gpp.21.919) also extend NS-management capabilities with admission-control policies, energy-efficiency indicators, and positioning/synchronization attributes, which support stronger SLA isolation and more sustainable operations. By codifying NRM objects, states, and notifications, these standards provide the foundation for E2E slice orchestration, dynamic adaptation, and exposure of NSaaS through defined open APIs. This framework remains consistent with the SBA vision of 5G and beyond while retaining flexibility for future enhancements. The instantiation of a Network Slice Instance (NSI) follows a standardized API sequence among Operations Support System (OSS)/ Business Support System (BSS), CSMF, NSMF, domain NSSMFs, NFVManagement and Orchestration (MANO), and the Inventory. Figure 3 illustrates the high-level flow and the interplay between GSMA, 3GPP/ETSI, and ETSI standards in this process. The process begins when the OSS/BSS captures a tenant request encoded with GSMA GST/NG.116, translating commercial requirements into a service order as specified by GSMA-NG116. The CSMF produces a ServiceProfile that maps high-level SLA attributes (throughput, latency, coverage, isolation) into a standards-compliant representation (3gpp.28.541).
Figure 3: End-to-End NS Management and Orchestration Across GSMA, 3GPP/ETSI, and ETSI Standards. Automation Scopes (AS1–AS4) follow the ETSI ZSM model for the evaluation, selection, configuration, and deployment of slice subnets.
The NSMF interprets this profile and issues allocateNsi, which creates context for the new slice and initiates decomposition into domain-specific subnets (3gpp.28.531). Each NG-RAN, TN, and 5GC NSSMF receives allocateNssi with a SliceProfile that refines resource constraints, including spectrum allocations, transport bandwidth, and core-function placement rules (3gpp.28.531). The NSSMF performs design-time resolution to select building blocks, policies, and placement options, producing a resolved subnet design aligned with declarative operations in TS 28.531 (3gpp.28.531). It engages the NFVMANO stack to instantiate the required Virtual Network Functions, Cloud-native Network Functions, and network resources so that performance objectives are considered from the outset (ETSI-NFV-MANO). As resources become active, the NSSMF invokes createMOI to persist the corresponding Managed Object Instances in the Inventory, formalizing their existence and relationships (3gpp.28.622). The NSMF consolidates these MOIs into an E2E slice, linking subnets and updating state to reflect readiness.
Finally, it is essential to highlight that where adjustments are required, modifyMOIAttributes is used to fine-tune runtime parameters such as slice capacity, policy bindings, or service exposure (3gpp.28.623). In this way, the Inventory evolves into a live repository of the slice structure, supporting lifecycle tracking and auditing. Closed-loop assurance systems and exposure APIs allow incremental modifications (via modifyMOIAttributes) that help maintain SLA compliance across NG-RAN, TN, and 5GC domains (etsi-zsm). Binding GST $\rightarrow$ ServiceProfile $\rightarrow$ SliceProfile $\rightarrow$ runtime MOIs preserves business intent and provides traceability from intent to realization. In accordance with the ETSI ZSM framework described by etsi-zsm, these steps correspond to Automation Scopes AS1–AS4, encompassing (AS1) candidate evaluation, (AS2) candidate selection, (AS3) aggregation and configuration of slicing entities, and (AS4) generation of subnet templates and operational deployment. Coordinated use of allocateNsi, allocateNssi, createMOI, and modifyMOIAttributes across all actors supports repeatable, interoperable, and automatable slice creation in line with ETSI ZSM’s zero-touch management as detailed by etsi-zsm.
3 Related Work
Table 1: Representative Related Work.
Works Type B5G/AI Ready Project Model Architecture Service Automation E2E Service LC Mgmt. NSaaS
ONAP Project L H H M M L
OSM Project L H H M M L
5GTours Project L H H M M L
5g-zorro Project M M M L L L
Baranda2020 Project L H H L L L
tranoris2021openslice Project M H H M M L
camara2025 Project M H H M
opencapif2025 Project M H M M L L
nephio2025 Project H H H H M M
sylva2024 Project M H H M L L
hexaX2 Project H M H H M M
monb5g Project H M M H M L
daemon Project H M H H M L
Li2018DeepRL Article L H L L L L
Jiang Article M H L H M L
Bega Article M H L L H L
Abbas2020IBNSlicingIN Article M H M H M L
Fernandez Article M H M M L L
Theodorou Article H H L H H L
Larrea2023 Article M H L L M L
Scotece2023-5gkube Article M M M H L L
mlopez2023mapek Article H L M H M L
Wyszkowski2024 Article M L H L L L
Dalgitsis2024-kp Article H H M M H L
Esmat2024 Article H H H M H L
Chowdhury2024 Article H H H H M L
mlops2024 Article H L H H H M
zhao2025 Article H H H H M L
NASP (This Work) Article H H H H H H
Maturity: Low (L), Medium (M), High (H); “–” = not applicable.
The recent literature on NS for 5G and beyond networks has centered on both E2E slice design and dynamic orchestration. Table 1 provides a qualitative comparison of the surveyed solutions. The ’Works’ column lists each solution, and Type distinguishes between runnable open-source projects and concept-focused articles. The B5G/AI Ready score reflects how well a study incorporates forward-looking Beyond Fifth Generation (B5G) concepts and AI-driven automation, from legacy 5G (L) to full AI/ML integration (H). Project Model rates the maturity of released software artifacts, ranging from a fully maintained public repository (H) through partial or prototype code (M) to no public code (L). Architecture captures the breadth of the E2E slice design across RAN, TN, and CN domains. Service Automation records adherence to ETSI ZSM closed-loop orchestration principles, and E2E Service LC Mgmt. measures the degree of automation across preparation, instantiation, assurance, and decommissioning. Finally, NSaaS indicates whether the work exposes NS as a tenant-facing service, progressing from no exposure (L), to domain-specific offerings (M), and, at the top end, full multi-domain NSaaS capability (H).
Several open-source initiatives automate 5G slice orchestration but vary in scope. ONAP and OSM offer comprehensive MANO frameworks but require significant integration effort for full E2E automation. Building on these, subsequent projects such as 5GTours, 5g-zorro, and Baranda2020 demonstrated vertical use cases but reported limited architectural evolution. Since 2023, the landscape has shifted toward B5G and AI-native solutions. Large-scale research programmes such as hexaX2 investigate B5G architectures that emphasize AI-driven automation and sustainability. Projects such as monb5g and daemon develop AI/ML-powered monitoring and control loops, aligning with the ETSI ZSM vision. The MAPE-K framework introduced by mlopez2023mapek is now a widely cited pattern for self-adaptive systems. Moreover, the MLOps practices cataloged by mlops2024 are being integrated into network automation to streamline the lifecycle of AI models in production networks. Furthermore, community-driven initiatives such as camara2025, opencapif2025, nephio2025, and sylva2024 develop open, cloud-native API and automation frameworks backed by major operators and vendors.
Pushing lifecycle automation further, OpenSlice, as described by tranoris2021openslice, adopts TM Forum’s Open Digital Architecture to drive catalog-based order decomposition that supports zero-touch onboarding, scaling, and retirement of NFV artifacts. However, it delegates granular 5GC tasks, such as UPF placement, PCF/ Network Repository Function (NRF) affinity, and network-exposure APIs, to external MANO stacks, excluding the SBA from the reconfiguration loop. In response, the NASP framework presented in this article unifies hierarchical slice composition, AI-powered assurance, and tenant-facing NSaaS exposure across RAN, TN, and CN domains.
AI-driven research is steadily closing specific gaps in the NS lifecycle. Li2018DeepRL were the first to apply Deep Reinforcement Learning (DRL) for joint radio- and core-resource balancing under fluctuating loads. Jiang introduced intelligence slicing, a unified framework that instantiates AI modules, such as Recurrent Neural Network (RNN)-based channel prediction and security-anomaly detection, on demand. Bega integrated a 3-D Convolution Neural Network (CNN) capacity-forecasting model (DeepCog) with Reinforcement Learning (RL) control loops, enabling proactive VNF scaling throughout the slice lifecycle. Abbas2020IBNSlicingIN proposed IBNSlicing, an intent-based, E2E framework that automatically translates high-level slice “contracts” into Open Source MANO (OSM) and FlexRAN templates; a hybrid Generative Adversarial Network (GAN), i.e., Long Short-term Memory (LSTM) + CNN, forecasts slice-level CPU and RAM requirements to trigger proactive scaling and admission decisions. Moreover, Fernandez designed a Distributed Ledger Technologies (DLT)-enabled 5G marketplace in which multiple providers securely advertise, negotiate, and trade heterogeneous resources via smart-contract automation, validating the concept with a Corda-based VNF-trading proof of concept. Complementing these point solutions, Wyszkowski2024 proposed a taxonomy of network-slice-instance patterns that now guides subsequent work.
Figure 4: NASP architecture.
Extending the state of the art, Theodorou outlined a blockchain-based, zero-touch service-assurance loop for cross-domain slicing, in which smart-contract-bound SLAs feed AI-driven monitoring, prediction, and automated reconfiguration across the lifecycle. Dalgitsis2024-kp proposed a cloud-native Slice Federation-as-a-Service (SFaaS) framework that augments the GSMA Operator Platform with a new east–west APIs, enabling seamless slice federation across operators and demonstrating the approach on a live 5G/B5G testbed. Esmat2024 decomposed carrier-grade SLAs into federated slice objectives, whereas Chowdhury2024 introduced Accelerator, an intent-driven, zero-touch resource-slicing algorithm that places and schedules Service Function Chain (SFC) workloads for B5G and legacy services across SDN/NFV infrastructures.
Cloud-native platforms are likewise paving the way for zero-touch cores. Scotece2023-5gkube presented 5G-Kube, a Kubernetes-native CNF stack that cuts slice instantiation times below 200 ms. Larrea2023 developed CoreKube, a cloud-native mobile core whose stateless, message-focused workers are containerized and orchestrated by Kubernetes. Additionally, zhao2025 combined adaptive RAN slicing with hierarchical DRL, achieving E2E KPIs across RAN, TN, and CN. Moreover, Wyszkowski2024 present a detailed tutorial that addresses a blind spot in slicing research and standardization: the design-time organization of slice and subnet instances. They systematize a taxonomy of slicing entities and formalize the slice/subnet design process, clarifying concepts not uniformly covered in 3GPP or ETSI. Building on this, the authors enumerate the essential information items a complete design must capture and propose a standards-aligned structure to record them. The tutorial decomposes design into fine-grained, ordered activities and offers automation guidelines that evolve from assisted handling to fully unattended, declarative workflows capable of dynamic slicing.
Despite these state of the art advances, three challenges remain: (i) E2E designs that span RAN, TN, and CN within a SBA; (ii) fully closed-loop slice lifecycle management aligned with SLA objectives; and (iii) an operational NSaaS framework that unifies these capabilities. The NASP framework addresses these gaps by (i) translating GSMA business templates into domain-specific descriptors, (ii) orchestrating hierarchical slice instantiation across RAN, TN, and CN via open interfaces, such as Radio Access Network Interface (RN1), Transport Network Interface (TN1), and Core Network Interface (CN1); and (iii) embedding AI-driven closed-loop assurance. Compared with other solutions, NASP integrates native B5G support, run-time optimization, and a demonstrable NSaaS prototype, illustrating a viable path toward zero-touch, ZSM-compliant B5G deployments.
4 NSaaS Platform Overview
This section introduces NASP, which aligns with concepts from 3GPP, ETSI, and GSMAframeworks to enable E2E slicing. The platform mediates between business intent and network implementation by automatically translating high‑level business templates into NS definitions, enabling multi‑domain resource orchestration and monitoring. In this context, NASP translates GSMA business templates into versioned Network Slice Templatess through a hierarchical deployment model. Business requests collected by the CSMF are transformed into service requirements and forwarded to the NSMF, which assigns slice identifiers (IDs) and delegates domain‑specific tasks. Standardized interfaces (RN1, TN1, and CN1) ensure seamless integration across the RAN, TN, and CN. Automation is achieved using CI/CD and IaC practices, while a closed‑loop monitoring system enforces continuous telemetry‑based SLA compliance.
Figure 4 presents the NASP reference architecture, comprising five main elements. The Onboard module (i) translates tenant requests into domain-specific requirements, leveraging GSMA and 3GPP standards, as well as operator-specific parameters, to ensure that high-level business templates are accurately mapped to the corresponding slice ( Network Slice Template (NST)/NSSTs) definitions. The Data Management module (ii) stores and manages all slice descriptors and artifacts through a domain-agnostic catalog and inventory, ensuring version control, lifecycle tracking, and consistent deployment across network domains. The Orchestration module (iii), fulfilling the roles of both CSMF and NSMF, coordinates slice instantiation across multi‑domain networks and incorporates functions such as data translation and resource mapping, as depicted by the integrated Resource Translations block. The Quality Assurance Closed Loop module (iv) monitors performance via closed‑loop telemetry and employs AI‑driven analytics to detect anomalies, interfacing with CI/CD tools to trigger automated reconfigurations that maintain SLA compliance. Finally, the Interfaces (v) RN1, TN1, and CN1 ensure connectivity among domain controllers, facilitating uniform management and synchronization across physical and virtual infrastructures.
Onboard module
This NASP module forms the northbound interface through which tenants submit slice requests encoded using the GSMA GST. Incoming service intents first pass through the Requirements micro‑service, where SLA parameters such as latency, throughput, and availability are normalized. The data stream is forwarded to Template Matching, whose rule‑based engine correlates the normalized attributes with a 3GPP‑compliant library of versioned NSSTs, selecting the candidate that minimizes resource footprint while maximizing reuse.
Once a suitable template is chosen, Resource Translations converts logical requirements into concrete descriptors for virtual compute, storage, radio, and transport resources, using mapping rules derived from TMF638; TMF639 and 3gpp.28.541. The translated artifact is forwarded to the Descriptions service, which renders a standards‑compliant Network Slice Descriptor (NSD) in YAML Ain’t Markup Language (YAML) format, directly consumable by the orchestration tier. A GitOps‑driven CI/CD pipeline performs syntax validation, policy checks, and version tagging before committing the artifact to an immutable registry. During onboarding, each GST attribute, such as e2eLatency, serviceAvailability, is persisted as a ServiceSpecification/SlaTarget object via TeleManagement Forum (TMF) 638 (v5.0.0). These targets are version‑controlled alongside the NS descriptors and immediately promoted to TMF 640 (v5.0.0) Service Level Objective (SLO) records so that downstream orchestration and assurance functions can reference a single, immutable source of truth.
Data Management module
Figure 5: Sequence diagram of slice resource allocation.
All slice descriptors and artifacts are stored in a domain-agnostic Catalog, which maintains versioned NSDs, NSSTs, virtual network function images, and Helm Chart Package Manager (Helm) charts. This repository provides template definitions for domain-specific orchestrators, ensuring that instantiation across domains follows consistent and deterministic patterns. A complementary Inventory does not duplicate these descriptors. Instead, it tracks the current lifecycle state of each slice instance across the underlying physical and virtual resources. It exposes authorised state information to admission control and assurance functions. Leveraging these resources, the Onboard module translates the tenant’s business intent into service definitions that can be processed by the orchestration tier without human intervention. The Catalog serves as the design-time repository for immutable templates, such as NSDs, NSSTs, Virtual Network Function Descriptors, and Helm charts, validated and version-controlled through CI/CD. The Inventory captures the runtime state of instantiated NSIs and Network Slice Subnet Instances, including lifecycle attributes, Key Performance Indicators, and SLA compliance. Overall, the Catalog defines what can be deployed, and the Inventory records what is deployed. This separation preserves traceability without duplication and supports federation through abstracted, authorized views.
Orchestration module
A four-tier hierarchy outlines responsibilities: Tier-1 (CSMF) captures business SLAs and converts them into service requirements; Tier-2 (NSMF) assigns slice identifiers and orchestrates tasks across domains; Tier-3 domain-specific NSSMFs, RAN, TN, and CN, allocate resources and instantiate slices from the descriptors stored in the Catalog; and Tier-4 comprises the actual physical or virtual instances configured from those descriptors. Rather than duplicating data-management functions, the orchestrator treats the Catalog solely as a source of template information and queries the Inventory for the live state of slice resources during provisioning and monitoring.
The end-to-end workflow, illustrated in Figure 5, begins when the tenant submits a slice request to the CSMF (01); the request is forwarded to the NSMF, which reserves an Single Network Slice Selection Assistance Information (S-NSSAI), retrieves the corresponding slice templates and engages the CN NSSMF to deploy the required CN VNFs under the supervision of the Network Function Virtualization Orchestrator (NFVO) Controller (02–04). Once an AMF endpoint is available, the RAN NSSMF instantiates the next-generation Node B (gNB) components through the RAN Controller so that they attach to the correct AMF (05–07). After both RAN and CN domains are operational, the TN NSSMF instructs the Wide Area Network (WAN) SDN Controller to establish transport connectivity, reports completion to the NSMF and triggers the final activation confirmation that is relayed to the tenant via the CSMF (08–10). All configuration artifacts are stored in the Inventory through idempotent transactions, guaranteeing that repeated executions converge to a consistent slice state, and the final slice configuration is preserved for subsequent monitoring and lifecycle management.
The transformation from high‑level business commitments into enforceable artifacts occurs in four logical phases. (i) Production: GST attributes are normalized and stored as TMF 638 (v5.0.0) SlaTarget objects, each of which is linked, via a mapping table inside the Resource Translations micro‑service, to the corresponding TMF 639 (v5.0.0) resource identifiers, e.g., transport bandwidth pools or radio‑spectrum slices. (ii) Decomposition: the CSMF converts these targets into TMF 640 (v5.0.0) SLO records and distributes them to the NSMF and the domain‑specific NSSMFs, which embed the derived KPI thresholds in their local controllers. (iii) Enforcement: during operation, the Quality‑Assurance loop continuously compares real‑time telemetry with the injected thresholds; any breach raises a TMF 642 (v5.0.0) notification that automatically triggers scaling, relocation, or healing workflows through TMF 645 (v5.0.0) Assurance APIs. (iv) Tenant visibility: current SLA status is exposed through a TMF 645 (v5.0.0) /assurance/v1/monitorings endpoint, and tenants may subscribe to NEF webhooks for proactive breach alerts.
The orchestration workflow follows the 3GPP management and service specifications, including TS 28.531 and TS 28.541 for slice lifecycle coordination, as well as TS 23.501 and TS 38.413 for inter-domain reference points. While 3GPP does not mandate a domain instantiation order, the NASP prototype deploys CNRANTN to mirror runtime dependencies. CN is instantiated first, so the AMF can expose its control- and user-plane endpoints: N2, i.e., NG Application Protocol (NGAP) over Stream Control Transmission Protocol (SCTP), for signalling and registration, and N3, i.e., GPRS Tunneling Protocol for 5G (GTP-5G) over User Datagram Protocol (UDP) for user-plane traffic. The AMF publishes these endpoints via its management API, and they are persisted in the Inventory with Internet Protocol (IP) addresses and ports. Next, the RAN orchestrator configures gNB components to complete NG Setup and register with the selected AMF over N2. Once CN and RAN are active, their artifacts, such as gNB IPs, Virtual Local Area Network (VLAN) IDs, and N3 forwarding parameters, are provided to the TN orchestrator, which programs slice-aware connectivity via the SDN controller, e.g., OpenFlow Protocol (OpenFlow) intents and VLAN-based paths, for both N2 and N3 traffic between RAN and CN subnets. This sequence establishes control-plane signalling before data-plane activation, aligns with the layering hierarchy, and remains compliant with the 3GPP slice-management framework and information models defined in TS 28.541.
Quality Assurance Closed Loop module
This module supervises every active NS through a non‑real‑time control loop that begins with Telemetry Discovery, which retrieves multilayer observability data (counters, alarms, traces, and logs) via the NEF, RAN monitoring interfaces, and SDN probes. Although transport-domain SDN telemetry (OpenFlow and sFlow Protocol (sFlow)/ IP Flow Information Export (IPFIX)) is not natively covered by 3GPP TS 28 5xx, a lightweight adaptor inside the Data Translation micro-service enriches every record with the S-NSSAI and converts raw counters, such as port-tx_bytes, queue delay, packet-drop events, into their 3GPP-compliant counterparts, defined in 3gpp.28.552. The normalized KPIs are pushed (i) towards the NWDAF for analytics or (ii) as TMF 642 events for fault management. This translation keeps the entire assurance loop 3GPP-compliant while still exploiting the high-resolution visibility provided by SDN probes. The records are normalized and time‑stamped before publication on a common bus. Data Aggregation aligns the heterogeneous streams on the slice ID, such as S‑NSSAI, and enriches them with business KPIs from the catalog. Finally, Data Translation harmonizes vendor‑specific semantics, converting resource counters into utilization metrics, mobility events into experience indicators, and sharing information into isolation scores, while mapping slice intents onto the configuration vocabulary of the Service Management and Orchestration (SMO), NSMF, and domain‑specific NSSMFs.
Service Quality Management can host a hierarchy of supervised and deep-RL models that at the decision tier can forecast QoS degradation, detect anomalies, and compute corrective actions such as scaling, function relocation, or policy updates, all guarded by SLA-aware safety checks. AI Mode & Training aims to orchestrate the learning lifecycle, retraining models on curated snapshots when concept drift appears and promoting validated artifacts to the online inference path without service disruption. Outcomes are disseminated through Events and Feeds, which stream real-time notifications to the operations dashboard, publish periodic assurance reports to tenants, and immutably log every action with before-and-after metrics to close the loop and enable continual improvement.
Interfaces
The NASP control plane exposes three domain‑agnostic southbound interfaces that enable uniform NS lifecycle automation across the RAN, TN, and CN domains. Each interface conveys an intent encoded as a YAML document that augments the 3GPP network‑resource model in TS 28.541 (v19.4.0, Rel‑19) with GST attributes defined by GSMA NG.116 (v10.0). RN1 targets the radio‑access network, TN1 the transport network, and CN1 the mobile core. For example, RN1 submits a RAN Slice Set to the Kubernetes (K8s) API server managing the cloud‑native RAN. The manifest is deployed through Helm releases, Network Attachment Definitions, and accompanying custom resources configuring the required gNB functions. Moreover, TN1 delivers a Path Intent over a RESTful Hypertext Transfer Protocol (HTTP)/2 channel to the SDN controller, translating the request into OpenFlow 1.5 flow‑mod messages. Finally, CN1 forwards a CN Slice Set to the K8s API server that orchestrates the CN. Therefore, the manifest is applied as Helm releases and ancillary resources, including a Data Plane Policy object that configures the user‑plane function datapath.
5 Prototype Implementation
This section presents the proof-of-concept built to validate the proposed NASP architecture for beyond-5G mobile networks. The prototype implements a context‑aware slice orchestrator in a containerized environment i.e., NASP’s Onboard, Orchestration, and Quality Assurance Closed Loop modules manage the end-to-end slice lifecycle (creation, update, and termination). It further extends the SMO concept by embedding business‑level objectives from GSMA templates directly into the orchestration workflow, bridging business intent and network service delivery. The prototype runs on Kubernetes v1.28 (Ubuntu 20.04 LTS) to manage containerized network functions across the RAN, TN, and CN domains. my5G-RANTester modules provide RAN functions, ONOS supervises a Mininet Virtual Network Emulator (Mininet) virtual environment to emulate the TN, and Free5gc implements the CN. The NASP prototype source code is available at: https://github.com/fhgrings/NASP.
The infrastructure uses Docker and the Kubernetes API server’s Representational State Transfer (REST) interface for configuration and management. The NASP control plane stores JSON-formatted NSSTs and NSTs in a document‑oriented database. Infrastructure descriptors are supplied through Helm charts and YAML manifests. The orchestrator applies context-aware logic that maps tenant metadata (e.g., latency budgets, UE density) to infrastructure-level configurations, dynamically adapting Helm chart parameters during instantiation. Monitoring, tracing, and alerting are provided by Prometheus, Istio, and Grafana, respectively. Control modules are written in Python 3.9; Flask handles HTTP requests, and Django serves a single‑page Web interface. Bash scripts coordinate low‑level Linux operations across domains.
Figure 6(a) depicts the NASP NST definition workflow, triggered when an operator submits a slice-template request. The orchestrator first checks for a tenant-provided definition; if present, the YAML/Helm fragment is parsed and validated. Otherwise, the engine derives the slice type (eMBB, URLLC, mMTC, or an operator-specific profile) from request metadata. The catalog is queried for versioned NSSTs per domain (RAN, TN, CN) together with version tags and resource footprints. For each exposed variable (e.g., maximum UE density, latency budget, sharing policy), the engine instantiates parameter sets and reserves artifacts (container images, Helm values, ConfigMaps). Finally, the selected NSSTs are composed into an E2E NST by generating a parent chart that imports the sub-charts and exports a consolidated values file. The resulting artifact is stored and marked as Ready for subsequent instantiation.
Figure 6(b) outlines the NSI instantiation workflow. When a customer requests an operational slice, the orchestrator generates a unique S-NSSAI, associates it with the tenant ID, and retrieves the corresponding NST produced in phase (a). Deployment proceeds domain-wise. First, CN NSSTs are instantiated to initialize control-plane functions and obtain the AMF access-point address. This address is then injected into the RAN slice template so that the gNB can bootstrap with the correct AMF target. Next, the orchestrator queries the inventory for RAN anchor points, updates ingress and egress intents on the TN controller, and deploys the TN NSST, which installs slice-specific forwarding rules. The workflow completes once both RAN and TN NSSTs report readiness, enabling UEs to attach and exchange data through the customized transport paths.
Figure 6: NASP workflows: (a) NST definition and (b) NSI instantiation.
NASP manipulates IP tables inside Kubernetes pods for TN configuration. Each pod’s network interfaces are dynamically reconfigured to route traffic using VLAN-based segmentation, providing isolation and meeting performance requirements (Figure 7). Distinct VLAN IDs enforce specific QoS policies: a short path prioritizes low latency, while a long path provides redundancy. Although the TN is programmed via ONOS/OpenFlow, we keep a lightweight 802.1Q tag on every packet. Distinct VLAN IDs encode slice-specific QoS policies: a short path prioritizes low latency for URLLC, whereas a longer path provides redundancy for mMTC and Shared slices. Although the TN is programmed via ONOS/OpenFlow, packets retain lightweight 802.1Q tags. These tags act as in-band slice and QoS markers, allowing switches to differentiate flows locally, reduce flow-table pressure, and perform deterministic path selection before any controller interaction. This mechanism preserves layer-2 isolation and remains compatible with legacy Ethernet segments (IEEE8021Q-2018).
Figure 7: Emulated TN topology with VLAN-based segmentation.
Implemented NASP Modules and Scope
Onboard module: Supports (i) intent ingestion via REST/JSON, (ii) slice-profile derivation from GSMA-aligned service templates (eMBB, URLLC, mMTC, non-3GPP), (iii) catalog lookup of versioned NSSTs, (iv) variable expansion and Helm-based template synthesis, and (v) persistence of rendered NSTs in a document-oriented store.
[…]
Quality Assurance Closed Loop module: Implements telemetry collection (Prometheus exporters), KPI extraction (deployment time, attach latency, availability, resource consumption), and threshold- or rule-based corrective actions such as AMF reconfiguration. Advanced AI-driven analytics (e.g., predictive scaling, anomaly classification) are currently placeholders for future work.
Interfaces: Implemented (RN1, TN1, CN1) as a unified northbound REST layer that normalizes slice intents into domain‑specific manifests (Helm/YAML for Kubernetes, JSON payloads for ONOS). Adaptors translate abstract bandwidth/latency/shareability fields into concrete Computer Processing Unit (CPU) quotas, pod replica counts, and flow‑rule priorities.
We consider Security and Policy Hooks as a basic enforcement mechanism via Kubernetes Role-Based Access Control (RBAC) and namespace scoping, plus slice-specific VLAN tagging and (for non-3GPP) Internet Protocol Security (IPsec) tunnel establishment through the Non-3GPP Interworking Function (N3IWF). Fine‑grained policy engines (AI‑driven admission) are future work.
Service Scenarios and Network-Function Mapping
We instantiated four slice scenarios to produce the results discussed later: (i) URLLC, (ii) mMTC (a.k.a., mIoT in later figures), (iii) Shared (an eMBB-type slice that reuses existing control‑plane Network Functions to measure best‑case provisioning), and (iv) non‑3GPP access (offload via N3IWF). Table 2 summarizes the architectural deltas that drive the different deployment and performance outcomes.
Table 2: Scenario‑specific mapping of 5G NFs, placement, and isolation mechanisms.
Scenario CN Functions RAN Config TN Path / VLAN Isolation Highlights
URLLC Dedicated AMF, SMF, UPF (edge); shared AUSF, UDR, NRF, PCF gNB CU/DU at edge; low PHY timer values Short 4-hop path; VLAN 101 (high priority) Dedicated control + data plane; CPU limits; exclusive flow rules
mMTC Shared AMF/SMF; lightweight dedicated UPF; enhanced UDR scaling Std. gNB; extended RA backoff (contention absorption) Long path (+2 hops resiliency); VLAN 102 Separate namespace; per-pod quotas; distinct S-NSSAI
Shared (eMBB) Reuses AMF/SMF/UPF; no new CN pods Std. gNB; default scheduler Same as mMTC path (reused flows); VLAN 102 Logical slice via S-NSSAI; policy QoS classes
non-3GPP Adds N3IWF + IPsec; dedicated UPF; shared AMF/SMF Minimal gNB role (Wi-Fi offload focus) Short path + tunnel endpoint; VLAN 104 IPsec tunnel; namespace + VLAN tag + S-NSSAI
Isolation Strategy
Isolation across the four scenarios is enforced in complementary layers. (1) Control‑plane logical isolation: Each slice obtains a unique S-NSSAI. Dedicated slices (URLLC) run separate AMF/ Session Management Function (SMF)/UPF pods in isolated Kubernetes namespaces, and shared slices rely on policy separation inside common pods. (2) Resource isolation: Kubernetes CPU/memory requests and limits reserve deterministic compute shares for dedicated NFs and HorizontalPodAutoscaler triggers (80% threshold) let URLLC replicas scale independently. (3) Data‑plane segmentation: VLAN tags (Table 2) plus ONOS-installed OpenFlow rules steer per‑slice traffic over distinct priority queues and (for URLLC) the minimal‑hop path. (4) Security/access isolation: non‑3GPP traffic traverses an IPsec tunnel terminated at the N3IWF and slice‑specific Kubernetes NetworkPolicies restrict pod reachability. (5) State isolation: Although AMF/SMF are stateless, per‑slice subscriber and session state is keyed by S-NSSAI in the Unified Data Repository (UDR) and cached via slice‑scoped ConfigMaps, preventing cross‑slice leakage. These details explain the performance differences observed later and provide the architectural context for the four evaluated service scenarios.
6 Evaluation Methodology
This section describes the methodology used to assess the performance of the proposed NASP architecture. The evaluation focuses on statistically robust and cost-aware metrics across five aspects: E2E slice-instantiation time, scalability, flexibility, cost efficiency, and UE connection latency. We perform the experiments on a cloud-native infrastructure that emulates realistic distributed topologies using public cloud resources, SDN WAN emulators, and containerized network function deployment.
Testbed Infrastructure and Network Topology
Experiments were executed on Virtual Machines provided by DigitalOcean. The testbed comprises heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0-88-generic. Control-plane components were deployed on s-4vcpu-8gb instances (4 vCPU, 8 GB RAM), whereas data-plane components used s-8vcpu-16gb instances (8 vCPU, 16 GB RAM). Dedicated VMs were allocated to Open Network Automation Platform (ONAP), Mininet with ONOS, and the Kubernetes cluster. The Kubernetes deployment (v1.28.3) used Calico (Calico) and Multus (Multus) to support multi-tenancy, with autoscaling thresholds configured at 80% for scale-up and 20% for scale-down. Persistent monitoring data were exported to a Prometheus–Grafana stack.
Experiments were conducted on VMs provided by DigitalOcean. The cloud testbed consists of heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0‑88‑generic. Control‑plane components ran on s‑4vcpu‑8gb instances (4 Virtual CPU (vCPU), 8 GB RAM), whereas data‑plane elements used s‑8vcpu‑16gb instances (8 vCPU, 16 GB RAM). We dedicated a separate VMs to ONAP, Mininet with ONOS, and a Kubernetes cluster. We used Kubernetes v1.28.3 with Calico and Multus to support multi‑tenancy, with autoscaling thresholds of 80% (scale‑up) and 20% (scale‑down). Persistent monitoring data were exported to a Prometheus–Grafana stack, and all timestamps were converted to Coordinated Universal Time (UTC) to avoid clock‑skew artifacts.
In this work, we propose three types of cloud sites for the network performance evaluation: Regional (Central), Metropolitan (Edge), and Internal (Extreme‑Low‑Latency Edge), as defined by gustavo2024evaluating. The Internal site offered limited instance types, so we normalized cost comparisons using pricing from all sites. We emulated the network topology using Mininet and the ONOS controller. Figure 8 shows the long‑range cloud‑network topology, where latency samples were collected from geographically separated regions. Metrics were sampled at 1 Hz over a 72-hour period. A cubic non-linear regression (adjusted $R^{2}_{\text{adj}}=0.93$) was used to summarize average-latency trends over a representative 24-hour cycle.
Figure 8: Long‑range cloud‑network topology.
Performance Metrics
The evaluation considers the following five aspects: E2E slice‑instantiation time (i): measured by decomposing design, configuration, and activation phases. Scalability (ii): assessed by increasing the number of slices and user connections without observable performance degradation (up to 250 concurrent UEs per slice). Flexibility (iii): inferred from the ability to customize slice requests and operate diverse 5G network functions profiles, including eMBB, URLLC, mMTC, and non‑3GPP. Cost efficiency (iv): quantified as the ratio between performance and operational expenses across heterogeneous instance types; we used the AWS TCO calculator to cross-check DigitalOcean pricing. UE connection latency (v): measured to capture the impact of network management and routing efficiency on end‑user experience. All latency statistics are reported as the median $\pm$ interquartile range (IQR) to mitigate the influence of outliers.
7 Experimental Evaluation
This section presents a comprehensive experimental evaluation of the proposed NASP solution. The results are organized along two dimensions: (i) the design and deployment of E2E NSIs and (ii) their performance and lifecycle management.
Figure 9: Slice deployment time across four scenarios (URLLC, mMTC, Shared eMBB reuse, non-3GPP offload).
The NASP design-and-deployment evaluation quantifies the time required for slice provisioning. Figure 9 shows the breakdown from the initial slice request to full deployment across the RAN, TN, and CN domains, including the subsequent auto-configuration phase. We compare four scenarios (Table 2): URLLC, mMTC, Shared (eMBB reuse), and non-3GPP offload. The Shared slice completes in approximately $22$ s because it reuses existing Core NFs (no new AMF/SMF/UPF instantiation). URLLC (${\approx}53$ s) and mMTC (${\approx}42$ s) require additional control-plane or stateful data-store preparation. The non-3GPP slice (${\approx}50$ s) adds N3IWF and IPsec setup overhead. These timing differences follow directly from the per-scenario NF placement, reuse, and isolation choices in Table 2, confirming that the design workflow (S- Network Slice Selection Assistance Information (NSSAI) assignment, NF selection/sizing, resource allocation) is efficient and scalable.
The slice-instantiation workflow is decomposed into N=26N{=}26 ordered steps and substeps (Figure 10). Step 1 creates the S-NSSAI identifier and triggers the initial CN deployment batch. Step 2 completes the CN control-plane instantiation. Step 3 covers the RAN domain, where only an emulated RAN is considered via my5G-RANTester. Step 4 schedules the eight full-duplex TN routes (four SDN hops $\times$ two directions) that connect the RAN and CN across the emulated WAN. Steps 5–6 execute NASP verification and issue the slice-activation handshake. We define per-step efficiency as the ratio between the number of provisioning operations executed within a step (e.g., Helm-chart releases, flow-rule insertions, database transactions) and the elapsed time of that step, expressed in actions $\text{s}^{-1}$. A larger value therefore indicates that the orchestrator completes more work per unit time, rather than merely the step being short. The step-wise breakdown suggests that reducing the number of NFs redeployed, particularly in Step 2, can lower the overall provisioning time. Step 2 corresponds to the CN phase in which CN NSSMF instantiates or reconfigures the control plane. Because this stage involves (i) pulling container images, (ii) applying Kubernetes manifests, and (iii) waiting for readiness probes across dependent microservices, it accounts for roughly 40 % of the total deployment time.
Figure 10: Per-step deployment time across network domains (CN, RAN, TN).
Figure 10 breaks down the contribution of each deployment step per domain. The CN domain exhibits the longest deployment times due to NF-to-NF internal communication and the overhead of Helm chart instantiation. The RAN domain exhibits intermediate delays, driven by the my5G-RANTester emulation components that initialize the logical gNB pods and execute streamlined configuration scripts. By contrast, the TN domain exhibits more agile behavior due to its direct interaction with its dedicated controller. This phase involves configuring the eight full-duplex TN routes (4 hops × 2 directions), i.e., all flow rules required by the emulated TN topology. These observations reinforce that streamlined control domains, such as TN, benefit from specialized controllers that minimize latency and improve overall efficiency of deployment.
The slice-lifecycle evaluation investigates UE attachment latency, dynamic slice reconfiguration, and resource utilization. Figure 11 shows the distribution of UE connection latencies in the four scenarios. These latency profiles are the result of the scenario-specific NF placement (Table 2), routing choices (Figure 7), and orchestrator sequencing. The URLLC scenario registers the lowest latency, due to the proximity of critical control functions and an exclusive TN route. The mMTC (mIoT) slice exhibits a moderate median with a pronounced tail caused by random-access congestion among many battery-constrained devices. By contrast, the Shared scenario incurs higher latency due to resource sharing among UEs.
The arrangement observed in Figure 11 occurs from four specific NASP design decisions: (i) URLLC keeps AMF/SMF replicas at the internal-edge site and installs a short, dedicated VLAN path (Figure 7), holding the control-plane Round-Trip Time (RTT) below 3 ms. (ii) mIoT uses the same edge placement, but random-access congestion in the RAN dominates join latency, creating the long tail. (iii) Shared multiplexes CN functions and TN routes across slices, so queueing inside common CN pods inflates the median attach time. (iv) Non-3GPP adds an extra N3IWF gateway, along with an IPsec tunnel setup (step 4 in Figure 10), which inserts roughly 1–1.5 s before AMF messages reach the CN. Moreover, the non-3GPP slice experiences additional attachment delay from IPsec tunnel establishment and interworking procedures required for offloading traffic from untrusted access. Values beyond $\pm 3\sigma$ were discarded as outliers to strengthen statistical significance and, hence, the reliability of the reported trends.
Figure 11: UE connection time across four scenarios (architectural mapping in Figure 4 and TN-route types in Figure 7).
Dynamic reconfiguration was characterized by monitoring slice availability (a binary metric, where 0 represents service outage and $1$ represents full operability) and UE registration latency during runtime updates of the CN and its supporting virtual infrastructure. The specific reorganization analyzed here transparently migrates the slice from an initial, low‑capacity control‑plane deployment to a higher‑capacity configuration that meets a sudden surge in QoS and throughput requirements while keeping ongoing sessions untouched. Figure 12 shows that NASP orchestration maintains service continuity even during the 9 s reconfiguration window delimited by the vertical dashed lines.
Figure 12: Slice availability during runtime AMF replace-and-redeploy procedure: the controller deletes the old AMF pod and launches a new one with an upgraded resource profile. The five vertical arrows indicate, in order, the start of Step 2.2, the outage interval while Step 2.2 executes, the end of Step 2.2, the completion of Steps 2.3–2.10, and the first successful UE registration in Step 6 (see Figure 10).
We also analyze UE-registration latency before, during, and after slice reconfiguration. Figure 13 shows the latency of the reconfiguration process with a 500 ms sampling period under NASP orchestration. In this run, the latency peaks at 1500 ms and quickly returns to the pre-reconfiguration baseline of approximately 600 ms. The peak reflects limited computing resources and communication across multiple networks and subnetworks in the evaluation scenario. The arrow labeled “Reconfig. start” (10 s) indicates when the controller begins destroying and redeploying the AMF pods. The subsequent arrow, “Timeout” (23 s), marks when the UE registration procedure stops receiving responses because no AMF instance is reachable. Service availability is restored at “NASP reconfig. end” (30 s) when the new AMF replica becomes ready. At 40 s, “Partial-dynamic reconfig. end” marks the completion of the partial dynamic orchestration, and “First response after reset” (47 s) highlights the first successful UE registration after the process. Destroying and redeploying the AMF pod is necessary because the new configuration, such as the higher replica count and tightened CPU quotas, modifies immutable Kubernetes fields. Therefore, Helm performs a recreate rather than an in-place update so that the fresh instance boots with the correct resource profile and slice identifiers. Since UE and session contexts reside in the external UDR, the AMF is stateless, and the brief $\approx 9$ s outage in Figure 12 does not affect the data plane. Moreover, Figure 13 shows the “No orchestration” latency. After reconfiguration starts at 10 s, latency rises sharply. No successful registration responses are observed between 23 s and 47 s, the exact interval highlighted by the “Timeout” arrow, because all attempts expire while the control plane is rebuilding.
Figure 13: Registration latency during the reconfiguration process.
We assessed the adaptability of slice management by examining slice‑transition events triggered when the number of concurrent UE connections exceeded a predefined threshold. Figure 14 shows that NASP triggers reconfiguration to maintain QoS without interruption. Until $t\simeq 51$ s, every new UE is admitted to “Slice with orchestration”, and its utilization grows linearly until it reaches the configured cap of seven simultaneous connections. At this instant, the NSSMF controller finalizes the reconfiguration (vertical dashed line) and redirects subsequent registration requests to “Slice without orchestration”. From that moment on, the green curve saturates while the blue curve rises, evidencing the controller’s ability to preserve the SLA of the saturated slice while still accommodating additional traffic. The different slopes of the curves follow directly from the admission‑control policy rather than any performance disparity between the slices. This dynamic switch indicates that the NASP architecture can support runtime scalability, ensuring that high-demand scenarios do not compromise overall network performance.
Figure 14: Slice utilization before and after dynamic reconfiguration (admission control event at dashed line).
We also monitored resource utilization continuously during slice deployment. Figure 15 presents combined vCPU and Random Access Memory (RAM) usage across the testbed, with averages of approximately 1.2 vCPU and 600 MB RAM over the NSI lifecycle. Monitoring these metrics highlights the efficiency of resource allocation and the small fluctuations resulting from automation-induced bursts, which are crucial for fine-tuning cost-management strategies in different deployment environments. The three vertical dash‑dotted arrows superimposed on Figure 15 mark resource demand: (i) the left-most arrow identifies when the batch of five slice requests is submitted, (ii) the central arrow marks the transition from request processing to slice instantiation, and (iii) the right-most arrow denotes when all five slices are fully deployed, at which point the infrastructure peaks at about 4.8 vCPU and 17.6 GB of RAM. These markers help relate each lifecycle stage to its corresponding CPU and memory consumption.
Figure 15: vCPU and RAM utilization over the NSI lifecycle.
Finally, we estimate the evaluation cost by relating the consumed computational resources to their operating costs in each environment (Edge, Metropolitan, Central (Cloud)). Unit prices were taken from public, on‑demand Amazon Web Services tariffs and summarized in Table 3. For each environment, we selected the closest general‑purpose flavor available in that tier. Monthly prices were normalized by advertised resources so that the regression coefficients express the marginal cost, in $\mathdollar\,\text{month}^{-1}$, of one vCPU and one GiB of RAM.
Table 3: Description of Edge, Metropolitan, and Central tier instances (amazonStartedWith).
Type Size vCPU RAM (GB) Storage (GB) Price/Month
Edge medium 2 4 200 \$70,88
Edge xlarge 4 16 200 \$193,52
Edge 2xlarge 8 64 200 \$526,40
Metropolitan medium 2 4 200 \$67,96
Metropolitan xlarge 4 16 200 \$117,60
Metropolitan 2xlarge 8 64 200 \$181,84
Central medium 2 4 200 \$46,37
Central xlarge 4 16 200 \$76,74
Central 2xlarge 8 64 200 \$137,47
We fit a linear regression to the values in Table 3, yielding one equation per environment: $\text{Edge}=39.42\,\text{CPU}+3.65\,\text{RAM}-22.56$, $\text{Metropolitan}=33.58\,\text{CPU}-1.46\,\text{RAM}+6.63$, and $\text{Central}=15.19\,\text{CPU}+1\,\text{RAM}+15.99$. We then used the consumption data from the previous evaluation to predict the price for each environment. As shown in Figure 16, the dash–dotted vertical lines mark (i) the submission of the first batch of five slice requests at $t=10$ s, (ii) the completion of that batch’s deployment at $t\approx 65$ s, (iii) the submission of a second batch of five slice requests at $t=60$ s, and (iv) the instant when all ten slices are fully operative at $t\approx 115$ s. Each marker corresponds to an inflection point in the Edge, Metropolitan, and Central cost curves, illustrating how incremental vCPU and RAM allocations immediately increase the predicted monthly costs.
Figure 16: Predicted monthly cost as a function of vCPU and RAM utilization.
Overall, these experimental results validate the capability of the NASP architecture to efficiently support network‑slice design, deployment, and dynamic lifecycle management. Beyond demonstrating flexibility across four 5G scenarios, the evaluation indicates a balanced trade‑off between deployment speed, run‑time performance, and operational cost. In this context, NASP provide a viable platform for next‑generation network slicing.
8 Conclusion
The evolution of 5G mobile networks is catalyzing digital transformation across many sectors, with standardization bodies such as ETSI and 3GPP, driving major architectural and technological advances. In this context, NSaaS is increasingly adopted as a foundation for flexible, tenant-oriented connectivity; however, its E2E realization still lacks a holistic standard. This work addresses this gap by proposing the NASP architecture, which maps business-level GSMA templates to technical NSI definitions and establishes connectivity across the RAN, TN, and CN domains. NASP is designed to interact with domain controllers, such as K8s and ONOS, and with standard Linux tools, such as iptables, namespaces, and Control Groups (CGROUPS), to enforce configuration changes, for example namespace routing and VLAN creation.
Evaluation across four scenarios demonstrates the platform’s capability to translate SLA specifications into E2E slice configurations, automate slice instantiation and management, and balance adaptability with resource consumption. Three quantitative findings are notable: (i) approximately 66% of the instantiation time is incurred in the CN domain, (ii) the URLLC slice achieves a 93% reduction in data-session establishment time compared with the Shared slice, and (iii) there is a 112% difference in monthly cost between Edge and Centralized deployments under identical performance targets.
Beyond the architectural contributions, the seamless integration of standardized GSMA templates, the unification of definitions across major mobile network entities to realize an E2E architecture, and the capability to instantiate network slices for Non-3GPP (non3GPP) applications open several avenues for future work. Planned efforts include: (a) integrating a physical RAN testbed to eliminate access-layer abstractions, (b) scaling the orchestration logic to support many concurrent slices, (c) embedding security and privacy assurance loops, (d) validating interoperability with commercial MANO stacks, such as ONAP and OSM, while aligning interfaces with emerging Open Radio Access Network (O-RAN) specifications in multi-operator environments, and (e) enhancing mobility management for seamless session continuity, particularly in light of the observed slice-establishment times.
Acknowledgment
This work was partially supported by CNPq Grants Nos. 405111/2021-5 and 130555/2019-3, and by CAPES, Finance Code 001, Brazil; additional support was provided by RNP and MCTIC under Grant No. 01245.010604/2020-14 as part of the 6G Brasil and OpenRAN@Brasil projects, and by MCTIC/CGI.br/FAPESP through Project SAMURAI (Grant No. 2020/05127-2) and Project PORVIR-5G (Grant No. 2020/05182-3); it has also been partially funded by the project XGM-AFCCT-2024-5-1-1 supported by xGMobile - EMBRAPII - Inatel Competence Center on 5G and 6G Networks, with financial resources from the PPI IoT/Manufatura 4.0/the MCTI grant number 052/2023, signed with EMBRAPII.

### v3

@@
 Results demonstrate the platform’s adaptability in handling diverse requests via the Communication Service Management Function.
-Evaluation indicates that Core configuration accounts for 68% of the time required to create a Network Slice Instance.
-Tests reveal a 93% reduction in data session establishment time when comparing URLLC and Shared scenarios.
+Evaluation indicates that the Core configuration accounts for 68% of the time required to create a Network Slice Instance.
+Tests reveal a 93% reduction in data session establishment time when comparing URLLC and Shared scenarios, i.e., an eMBB-type slice that reuses existing control-plane Network Functions to represent a best-effort provisioning baseline.
 Finally, we present cost variations for operating the platform with the orchestration of five and ten slices, showing a 112% variation between Edge and Central deployments.
@@
 1 Introduction
-The rapid, sustained increase in mobile data traffic, combined with the emerging diversity of service requirements from Massive Machine-Type Communication (mMTC) terminals and vertical enterprises, challenges the traditional mobile‐broadband paradigm (Navarro-Ortiz2020).
-Legacy networks, initially designed with only a limited set of parameters (e.g., priority class and Quality of Service (QoS) for broadband services), now face the task of accommodating heterogeneous and, in many cases, conflicting operational needs (Vila2020).
+The rapid, sustained increase in mobile data traffic, combined with the emerging diversity of service requirements from Massive Machine-Type Communication (mMTC) terminals and vertical enterprises, challenges the traditional mobile‐broadband paradigm (Navarro-Ortiz and others, 2020).
+Legacy networks, initially designed with only a limited set of parameters (e.g., priority class and Quality of Service (QoS) for broadband services), now face the task of accommodating heterogeneous and, in many cases, conflicting operational needs (Vilà and others, 2020).
 For instance, a railway network demands seamless, high-speed mobility management along predefined routes, whereas an electricity metering system requires support for static, low-volume transmissions.
-Recent advances in Network Functions Virtualization (NFV) and Software-Defined Networking (SDN) have paved the way for innovations such as Network Slicing (NS), which enables the logical separation of network functions and resources tailored to specific technical and commercial requirements (Wijethilaka2021; Ksentini2020; Mahdi2023).
-NS has conceptual roots in earlier technologies, such as the 802.1Q virtual local‑area network standard (IEEE8021Q-2018) and Internet Engineering Task Force (IETF) RFC 4026 Virtual Private Networks (rfc4026), which provided isolated broadcast and session domains.
-However, deploying NS in cellular environments introduces additional challenges, including mobility management, control‑plane authentication, and session and charging management in the user plane (mahyoub2024security).
-Studies led by the Third Generation Partnership Project (3GPP) and initiatives by major industry players indicate an evolution toward End-to-End (E2E) network slices that integrate both 3GPP and non‑3GPP domains, enhancing connectivity for the Internet of Things (IoT) and enabling diverse service customization (Nguyen2016; bertenyi2018ng; Linnartz2022-lb).
-Despite these advancements, the complexity stemming from distributed interfaces across Radio Access Network (RAN), Transport Network (TN), and Core Network (CN) necessitates novel orchestration and management solutions (Wyszkowski2024; 7962822).
+Recent advances in Network Functions Virtualization (NFV) and Software-Defined Networking (SDN) have paved the way for innovations such as Network Slicing (NS), which enables the logical separation of network functions and resources tailored to specific technical and commercial requirements (Wijethilaka and Liyanage, 2021; Ksentini and Frangoudis, 2020; Mahdi and Abdullah, 2023).
+NS has conceptual roots in earlier technologies, such as the 802.1Q virtual local‑area network standard (IEEE, 2018) and Internet Engineering Task Force (IETF) RFC 4026 Virtual Private Networks (Andersson and Madsen, 2005), which provided isolated broadcast and session domains.
+However, deploying NS in cellular environments introduces additional challenges, including mobility management, control‑plane authentication, and session and charging management in the user plane (Mahyoub and others, 2024).
+Studies led by the Third Generation Partnership Project (3GPP) and initiatives by major industry players indicate an evolution toward End-to-End (E2E) network slices that integrate both 3GPP and non‑3GPP domains, enhancing connectivity for the Internet of Things (IoT) and enabling diverse service customization (Nguyen and others, 2016; Bertenyi and others, 2018; Linnartz and others, 2022).
+Despite these advancements, the complexity stemming from distributed interfaces across Radio Access Network (RAN), Transport Network (TN), and Core Network (CN) necessitates novel orchestration and management solutions (Wyszkowski and others, 2024; Devlic and others, 2017).
 The evolution toward heterogeneous mobile networks, driven by massive IoT deployments, emerging business-oriented services, and stringent performance requirements, motivates the need for a comprehensive, automated approach to NS.
-Existing studies have addressed resource allocation using machine‑learning techniques (Jiang; Li2018DeepRL) and have developed frameworks for NS control and assurance (Abbas2020IBNSlicingIN; Theodorou; Bega).
+Existing studies have addressed resource allocation using machine‑learning techniques (Jiang et al., 2019; Li et al., 2018) and have developed frameworks for NS control and assurance (Abbas and others, 2020; Theodorou and others, 2021; Bega and others, 2020).
 However, fully integrated E2E orchestration that translates Service Level Agreement (SLA)/ Service Level Specification (SLS) requirements into practical NS designs remains an open challenge, as clear guidelines that bridge the gap between business-level intents and concrete slice configurations spanning the RAN, TN, and CN domains are still lacking in the literature.
@@
 Second, we propose a hierarchical orchestration model, including Communication Service Management Function (CSMF), Network Slice Management Function (NSMF), and domain‐specific Network Slice Subnet Management Functions, which integrate closed‐loop quality assurance via CI/CD and IaC practices, enabling dynamic SLA compliance through AI‐driven telemetry analytics.
-Third, we develop and release a fully functional prototype built on Kubernetes, Open Network Operating System (ONOS), my5G-RANTester (Silveira2022), and Free 5G Core (Free5gc), demonstrating seamless integration, extensibility, and native support for physical and virtual resources in heterogeneous deployment environments.
+Third, we develop and release a fully functional prototype built on Kubernetes, Open Network Operating System (ONOS), my5G-RANTester (Silveira and others, 2022), and Free 5G Core (Free5gc), demonstrating seamless integration, extensibility, and native support for physical and virtual resources in heterogeneous deployment environments.
 Our comprehensive evaluation, covering mMTC, Ultra-Reliable Low Latency Communications (URLLC), shared, and non-3GPP use cases, yields several impactful findings.
+In this context, the shared use case corresponds to an Enhanced Mobile Broadband (eMBB)-type slice that reuses existing control-plane Network Functions, serving as a best-effort baseline for comparison against fully isolated deployments.
 We observe that 66% of the E2E slice instantiation time is spent on Core domain configuration, pinpointing critical optimization targets.
@@
 2 Background and Evolution of 5G Networks
-The evolution of mobile telecommunications toward Fifth Generation (5G) introduced network programmability and NSaaS to better meet emerging service and market needs (3gpp.21.915; afolabi2018network).
-According to 3gpp.23.501, the 5G system and its Service-Based Architecture (SBA) were first specified in Releases 15–16, with continued refinements in later 3GPP releases (TS 23.501 v19.2.0, Rel-19).
+The evolution of mobile telecommunications toward Fifth Generation (5G) introduced network programmability and NSaaS to better meet emerging service and market needs (3GPP, 2019a; Afolabi and others, 2018).
+According to 3GPP (2025e), the 5G system and its Service-Based Architecture (SBA) were first specified in Releases 15–16, with continued refinements in later 3GPP releases (TS 23.501 v19.2.0, Rel-19).
 The architecture retains core elements, User Equipment (UE), Next Generation Radio Access Network (NG-RAN), and 5G Core (5GC), and separates control and user planes via Access and Mobility Management Function (AMF) and User Plane Function (UPF).
@@
 NS is a central 5G innovation, allowing multiple logical networks to share a common infrastructure while meeting distinct SLAs.
-Each slice orchestrates resources across the RAN, TN, and CN and is characterized according to 3GPP and Global System for Mobile Communications Association (GSMA) guidance (3gpp.23.501; gsma-116) (GSMA NG.116 v10.0).
+Each slice orchestrates resources across the RAN, TN, and CN and is characterized according to 3GPP and Global System for Mobile Communications Association (GSMA) guidance (3GPP, 2025e; GSMA, 2025) (GSMA NG.116 v10.0).
 The Network Slice Type (NEST) concept groups key attributes, throughput, latency, and reliability, into a service profile.
-For example, an Enhanced Mobile Broadband (eMBB) slice may target 99.999% availability, multimedia telephony, specific QoS, and strong session continuity.
-Moreover, isolation is essential since congestion or scaling in one slice must not degrade others, and separation must hold for performance, management, and security/privacy (3gpp.28.552).
+For example, an eMBB slice may target 99.999% availability, multimedia telephony, specific QoS, and strong session continuity.
+Moreover, isolation is essential since congestion or scaling in one slice must not degrade others, and separation must hold for performance, management, and security/privacy (3GPP, 2025d).
 In this case, the system can be viewed in three domains, as can be seen in Figure 1: the RAN (gNB with Radio Unit (RU)/ Distributed Unit (DU)/ Centralized Unit (CU)), the TN (access–core transport), and the CN (interfacing with external Data Networks via the UPF).
-Releases 17–18 (3gpp.21.917; 3gpp.21.918) extend slicing toward NSaaS, enabling operators to offer tenants customized, on-demand slices.
-NSaaS exposes management capabilities via standardized Application Programming Interfaces from 3GPP and TM Forum frameworks (gsma-116).
+Releases 17–18 (3GPP, 2022b, 2024c) extend slicing toward NSaaS, enabling operators to offer tenants customized, on-demand slices.
+NSaaS exposes management capabilities via standardized Application Programming Interfaces from 3GPP and TM Forum frameworks (GSMA, 2025).
 Slice management relies on CSMF, NSMF, and NSSMF to translate tenant demands into an E2E deployment.
@@
 Figure 2: NS orchestration architecture.
-NS lifecycle management, from preparation (blueprints, reservations) to operation (activation, monitoring, capacity planning) and decommissioning, benefits from modularity and isolation (5gcorebook-ns).
-The resulting operational complexity motivates Zero-touch Network and Service Management (ZSM), which applies automation, Machine Learning (ML), and Artificial Intelligence (AI) for self-configuration, self-monitoring, self-healing, and self-optimization (etsi-zsm).
+NS lifecycle management, from preparation (blueprints, reservations) to operation (activation, monitoring, capacity planning) and decommissioning, benefits from modularity and isolation (Rommer and others, 2020).
+The resulting operational complexity motivates Zero-touch Network and Service Management (ZSM), which applies automation, Machine Learning (ML), and Artificial Intelligence (AI) for self-configuration, self-monitoring, self-healing, and self-optimization (ETSI, 2024).
 Therefore, beyond-5G systems build on these principles to deliver agile, customized services that meet diverse SLAs.
-Building on these advances, 3gpp.21.917; 3gpp.21.918 describe the latest 3GPP specifications (TS 28.541, TS 28.531, TS 28.622, TS 28.623; Releases 18–19), and together with etsi-zsm they outline a precise Network Resource Model (NRM) for standardized management of 5GC, NG-RAN, and NS.
-Function Information Object Classes, such as AMFFunction, SMFFunction, UPFFunction, GNBDUFunction, GNBCUCPFunction, encapsulates logical behaviors with well-defined attributes, states, and relationships.
-The subnetwork IOC provides a hierarchical container to organize ManagedElements into administrative domains.
-For slicing, NetworkSlice and NetworkSliceSubnet IOCs are augmented with ServiceProfile and SliceProfile data types that represent performance, coverage, capacity, and isolation requirements, aligning with GSMA Generic Network Slice Template (GST)/NG.116 service descriptors to maintain globally consistent SLA representations.
-These definitions are bound to solution sets in Extensible Markup Language (XML), JavaScript Object Notation (JSON), and YANG, supporting interoperability across multi-vendor environments and aligning with European Telecommunications Standards Institute (ETSI) ZSM principles of zero-touch automation.
-Releases 18–19 (3gpp.21.918; 3gpp.21.919) also extend NS-management capabilities with admission-control policies, energy-efficiency indicators, and positioning/synchronization attributes, which support stronger SLA isolation and more sustainable operations.
-By codifying NRM objects, states, and notifications, these standards provide the foundation for E2E slice orchestration, dynamic adaptation, and exposure of NSaaS through defined open APIs.
-This framework remains consistent with the SBA vision of 5G and beyond while retaining flexibility for future enhancements.
+Building on these advances, the latest 3GPP specifications (Rel-18/19) outline a precise Network Resource Model (NRM) utilizing Information Object Classes (e.g., NetworkSlice, NetworkSliceSubnet) to standardize management (3GPP, 2024c).
+These are augmented with ServiceProfile and SliceProfile descriptors that align with GSMA Generic Network Slice Template (GST) to maintain globally consistent SLA representations.
+By codifying NRM objects, states, and operations across multi-vendor definitions ( Extensible Markup Language (XML), JavaScript Object Notation (JSON), YANG), these standards provide the necessary foundation for E2E slice orchestration and exposure of NSaaS through open APIs.
 The instantiation of a Network Slice Instance (NSI) follows a standardized API sequence among Operations Support System (OSS)/ Business Support System (BSS), CSMF, NSMF, domain NSSMFs, NFVManagement and Orchestration (MANO), and the Inventory.
-Figure 3 illustrates the high-level flow and the interplay between GSMA, 3GPP/ETSI, and ETSI standards in this process.
-The process begins when the OSS/BSS captures a tenant request encoded with GSMA GST/NG.116, translating commercial requirements into a service order as specified by GSMA-NG116.
-The CSMF produces a ServiceProfile that maps high-level SLA attributes (throughput, latency, coverage, isolation) into a standards-compliant representation (3gpp.28.541).
+Figure 3 illustrates the high-level flow and the interplay between GSMA, 3GPP/ European Telecommunications Standards Institute (ETSI), and ETSI standards in this process.
+The process begins when the OSS/BSS captures a tenant request encoded with GSMA GST/NG.116, translating commercial requirements into a service order as specified by GSMA (2023).
+The CSMF produces a ServiceProfile that maps high-level SLA attributes (throughput, latency, coverage, isolation) into a standards-compliant representation (3GPP, 2025a).
 Figure 3: End-to-End NS Management and Orchestration Across GSMA, 3GPP/ETSI, and ETSI Standards.
 Automation Scopes (AS1–AS4) follow the ETSI ZSM model for the evaluation, selection, configuration, and deployment of slice subnets.
-The NSMF interprets this profile and issues allocateNsi, which creates context for the new slice and initiates decomposition into domain-specific subnets (3gpp.28.531).
-Each NG-RAN, TN, and 5GC NSSMF receives allocateNssi with a SliceProfile that refines resource constraints, including spectrum allocations, transport bandwidth, and core-function placement rules (3gpp.28.531).
-The NSSMF performs design-time resolution to select building blocks, policies, and placement options, producing a resolved subnet design aligned with declarative operations in TS 28.531 (3gpp.28.531).
-It engages the NFVMANO stack to instantiate the required Virtual Network Functions, Cloud-native Network Functions, and network resources so that performance objectives are considered from the outset (ETSI-NFV-MANO).
-As resources become active, the NSSMF invokes createMOI to persist the corresponding Managed Object Instances in the Inventory, formalizing their existence and relationships (3gpp.28.622).
-The NSMF consolidates these MOIs into an E2E slice, linking subnets and updating state to reflect readiness.
-Finally, it is essential to highlight that where adjustments are required, modifyMOIAttributes is used to fine-tune runtime parameters such as slice capacity, policy bindings, or service exposure (3gpp.28.623).
-In this way, the Inventory evolves into a live repository of the slice structure, supporting lifecycle tracking and auditing.
-Closed-loop assurance systems and exposure APIs allow incremental modifications (via modifyMOIAttributes) that help maintain SLA compliance across NG-RAN, TN, and 5GC domains (etsi-zsm).
-Binding GST $\rightarrow$ ServiceProfile $\rightarrow$ SliceProfile $\rightarrow$ runtime MOIs preserves business intent and provides traceability from intent to realization.
-In accordance with the ETSI ZSM framework described by etsi-zsm, these steps correspond to Automation Scopes AS1–AS4, encompassing (AS1) candidate evaluation, (AS2) candidate selection, (AS3) aggregation and configuration of slicing entities, and (AS4) generation of subnet templates and operational deployment.
-Coordinated use of allocateNsi, allocateNssi, createMOI, and modifyMOIAttributes across all actors supports repeatable, interoperable, and automatable slice creation in line with ETSI ZSM’s zero-touch management as detailed by etsi-zsm.
+Following GST modeling, the NSMF translates the ServiceProfile into an allocateNsi request to decompose the slice into domain-specific subnets.
+Each domain NSSMF resolves the SliceProfile and engages the NFVMANO stack to instantiate resources, which are persisted as Managed Object Instances via createMOI 3GPP (2025b); ETSI (2021).
+When dynamic adjustments are necessary, runtime parameter tuning is achieved through modifyMOIAttributes (3GPP, 2025c).
+This structured progression from business intent (GST $\rightarrow$ ServiceProfile $\rightarrow$ SliceProfile) to runtime MOIs corresponds directly to the ETSI ZSM Automation Scopes AS1–AS4 (ETSI, 2024), enabling a repeatable, zero-touch orchestration process.
 3 Related Work
@@
 Works Type B5G/AI Ready Project Model Architecture Service Automation E2E Service LC Mgmt. NSaaS
-ONAP Project L H H M M L
-OSM Project L H H M M L
-5GTours Project L H H M M L
-5g-zorro Project M M M L L L
-Baranda2020 Project L H H L L L
-tranoris2021openslice Project M H H M M L
-camara2025 Project M H H M
-opencapif2025 Project M H M M L L
-nephio2025 Project H H H H M M
-sylva2024 Project M H H M L L
-hexaX2 Project H M H H M M
-monb5g Project H M M H M L
-daemon Project H M H H M L
-Li2018DeepRL Article L H L L L L
-Jiang Article M H L H M L
-Bega Article M H L L H L
-Abbas2020IBNSlicingIN Article M H M H M L
-Fernandez Article M H M M L L
-Theodorou Article H H L H H L
-Larrea2023 Article M H L L M L
-Scotece2023-5gkube Article M M M H L L
-mlopez2023mapek Article H L M H M L
-Wyszkowski2024 Article M L H L L L
-Dalgitsis2024-kp Article H H M M H L
-Esmat2024 Article H H H M H L
-Chowdhury2024 Article H H H H M L
-mlops2024 Article H L H H H M
-zhao2025 Article H H H H M L
-NASP (This Work) Article H H H H H H
+ONAP (2022) Project L H H M M L
+ETSI MANO (2022) Project L H H M M L
+5G Tours (2024) Project L H H M M L
+Breitgand and others (2021) Project M M M L L L
+Baranda and others (2020) Project L H H L L L
+Tranoris (2021) Project M H H M M L
+CAMARA Project Consortium (2025) Project M H H M
+ETSI TeraFlowSDN and OpenCAPIF Team (2025) Project M H M M L L
+Nephio Project Community (2025) Project H H H H M M
+Project Sylva Consortium (2024) Project M H H M L L
+Hexa-X-II Consortium / SNS JU (2025) Project H M H H M M
+MonB5G Project Consortium (2023) Project H M M H M L
+DAEMON Project Consortium (2023) Project H M H H M L
+Li et al. (2018) Article L H L L L L
+Jiang et al. (2019) Article M H L H M L
+Bega and others (2020) Article M H L L H L
+Abbas and others (2020) Article M H M H M L
+Fernández-Fernández and others (2021) Article M H M M L L
+Theodorou and others (2021) Article H H L H H L
+Larrea et al. (2023) Article M H L L M L
+Scotece and others (2023) Article M M M H L L
+López and others (2023) Article H L M H M L
+Wyszkowski and others (2024) Article M L H L L L
+Dalgitsis and others (2024) Article H H M M H L
+Esmat and Lorenzo (2024) Article H H H M H L
+Chowdhury (2024) Article H H H H M L
+Smith and Doe (2024) Article H L H H H M
+Zhao et al. (2025) Article H H H H M L
+NASP (This Work) Article M H H H H H
 Maturity: Low (L), Medium (M), High (H); “–” = not applicable.
@@
 Several open-source initiatives automate 5G slice orchestration but vary in scope.
-ONAP and OSM offer comprehensive MANO frameworks but require significant integration effort for full E2E automation.
-Building on these, subsequent projects such as 5GTours, 5g-zorro, and Baranda2020 demonstrated vertical use cases but reported limited architectural evolution.
+ONAP (2022) and ETSI MANO (2022) offer comprehensive MANO frameworks but require significant integration effort for full E2E automation.
+Building on these, subsequent projects such as 5G Tours (2024), Breitgand and others (2021), and Baranda and others (2020) demonstrated vertical use cases but reported limited architectural evolution.
 Since 2023, the landscape has shifted toward B5G and AI-native solutions.
-Large-scale research programmes such as hexaX2 investigate B5G architectures that emphasize AI-driven automation and sustainability.
-Projects such as monb5g and daemon develop AI/ML-powered monitoring and control loops, aligning with the ETSI ZSM vision.
-The MAPE-K framework introduced by mlopez2023mapek is now a widely cited pattern for self-adaptive systems.
-Moreover, the MLOps practices cataloged by mlops2024 are being integrated into network automation to streamline the lifecycle of AI models in production networks.
-Furthermore, community-driven initiatives such as camara2025, opencapif2025, nephio2025, and sylva2024 develop open, cloud-native API and automation frameworks backed by major operators and vendors.
-Pushing lifecycle automation further, OpenSlice, as described by tranoris2021openslice, adopts TM Forum’s Open Digital Architecture to drive catalog-based order decomposition that supports zero-touch onboarding, scaling, and retirement of NFV artifacts.
+Large-scale research programmes such as Hexa-X-II Consortium / SNS JU (2025) investigate B5G architectures that emphasize AI-driven automation and sustainability.
+Projects such as MonB5G Project Consortium (2023) and DAEMON Project Consortium (2023) develop AI/ML-powered monitoring and control loops, aligning with the ETSI ZSM vision.
+The MAPE-K framework introduced by López and others (2023) is now a widely cited pattern for self-adaptive systems.
+Moreover, the MLOps practices cataloged by Smith and Doe (2024) are being integrated into network automation to streamline the lifecycle of AI models in production networks.
+Furthermore, community-driven initiatives such as CAMARA Project Consortium (2025), ETSI TeraFlowSDN and OpenCAPIF Team (2025), Nephio Project Community (2025), and Project Sylva Consortium (2024) develop open, cloud-native API and automation frameworks backed by major operators and vendors.
+Pushing lifecycle automation further, OpenSlice, as described by Tranoris (2021), adopts TM Forum’s Open Digital Architecture to drive catalog-based order decomposition that supports zero-touch onboarding, scaling, and retirement of NFV artifacts.
 However, it delegates granular 5GC tasks, such as UPF placement, PCF/ Network Repository Function (NRF) affinity, and network-exposure APIs, to external MANO stacks, excluding the SBA from the reconfiguration loop.
@@
 AI-driven research is steadily closing specific gaps in the NS lifecycle.
-Li2018DeepRL were the first to apply Deep Reinforcement Learning (DRL) for joint radio- and core-resource balancing under fluctuating loads.
-Jiang introduced intelligence slicing, a unified framework that instantiates AI modules, such as Recurrent Neural Network (RNN)-based channel prediction and security-anomaly detection, on demand.
-Bega integrated a 3-D Convolution Neural Network (CNN) capacity-forecasting model (DeepCog) with Reinforcement Learning (RL) control loops, enabling proactive VNF scaling throughout the slice lifecycle.
-Abbas2020IBNSlicingIN proposed IBNSlicing, an intent-based, E2E framework that automatically translates high-level slice “contracts” into Open Source MANO (OSM) and FlexRAN templates; a hybrid Generative Adversarial Network (GAN), i.e., Long Short-term Memory (LSTM) + CNN, forecasts slice-level CPU and RAM requirements to trigger proactive scaling and admission decisions.
-Moreover, Fernandez designed a Distributed Ledger Technologies (DLT)-enabled 5G marketplace in which multiple providers securely advertise, negotiate, and trade heterogeneous resources via smart-contract automation, validating the concept with a Corda-based VNF-trading proof of concept.
-Complementing these point solutions, Wyszkowski2024 proposed a taxonomy of network-slice-instance patterns that now guides subsequent work.
+Li et al.
+(2018) were the first to apply Deep Reinforcement Learning (DRL) for joint radio- and core-resource balancing under fluctuating loads.
+Jiang et al.
+(2019) introduced intelligence slicing, a unified framework that instantiates AI modules, such as Recurrent Neural Network (RNN)-based channel prediction and security-anomaly detection, on demand.
+Bega and others (2020) integrated a 3-D Convolution Neural Network (CNN) capacity-forecasting model (DeepCog) with Reinforcement Learning (RL) control loops, enabling proactive Virtual Network Function (VNF) scaling throughout the slice lifecycle.
+Abbas and others (2020) proposed IBNSlicing, an intent-based, E2E framework that automatically translates high-level slice “contracts” into Open Source MANO (OSM) and FlexRAN templates; a hybrid Generative Adversarial Network (GAN), i.e., Long Short-term Memory (LSTM) + CNN, forecasts slice-level CPU and RAM requirements to trigger proactive scaling and admission decisions.
+Moreover, Fernández-Fernández and others (2021) designed a Distributed Ledger Technologies (DLT)-enabled 5G marketplace in which multiple providers securely advertise, negotiate, and trade heterogeneous resources via smart-contract automation, validating the concept with a Corda-based VNF-trading proof of concept.
+Complementing these point solutions, Wyszkowski and others (2024) proposed a taxonomy of network-slice-instance patterns that now guides subsequent work.
 Figure 4: NASP architecture.
-Extending the state of the art, Theodorou outlined a blockchain-based, zero-touch service-assurance loop for cross-domain slicing, in which smart-contract-bound SLAs feed AI-driven monitoring, prediction, and automated reconfiguration across the lifecycle.
-Dalgitsis2024-kp proposed a cloud-native Slice Federation-as-a-Service (SFaaS) framework that augments the GSMA Operator Platform with a new east–west APIs, enabling seamless slice federation across operators and demonstrating the approach on a live 5G/B5G testbed.
-Esmat2024 decomposed carrier-grade SLAs into federated slice objectives, whereas Chowdhury2024 introduced Accelerator, an intent-driven, zero-touch resource-slicing algorithm that places and schedules Service Function Chain (SFC) workloads for B5G and legacy services across SDN/NFV infrastructures.
+Extending the state of the art, Theodorou and others (2021) outlined a blockchain-based, zero-touch service-assurance loop for cross-domain slicing, in which smart-contract-bound SLAs feed AI-driven monitoring, prediction, and automated reconfiguration across the lifecycle.
+Dalgitsis and others (2024) proposed a cloud-native Slice Federation-as-a-Service (SFaaS) framework that augments the GSMA Operator Platform with a new east–west APIs, enabling seamless slice federation across operators and demonstrating the approach on a live 5G/B5G testbed.
+Esmat and Lorenzo (2024) decomposed carrier-grade SLAs into federated slice objectives, whereas Chowdhury (2024) introduced Accelerator, an intent-driven, zero-touch resource-slicing algorithm that places and schedules Service Function Chain (SFC) workloads for B5G and legacy services across SDN/NFV infrastructures.
 Cloud-native platforms are likewise paving the way for zero-touch cores.
-Scotece2023-5gkube presented 5G-Kube, a Kubernetes-native CNF stack that cuts slice instantiation times below 200 ms.
-Larrea2023 developed CoreKube, a cloud-native mobile core whose stateless, message-focused workers are containerized and orchestrated by Kubernetes.
-Additionally, zhao2025 combined adaptive RAN slicing with hierarchical DRL, achieving E2E KPIs across RAN, TN, and CN.
-Moreover, Wyszkowski2024 present a detailed tutorial that addresses a blind spot in slicing research and standardization: the design-time organization of slice and subnet instances.
+Scotece and others (2023) presented 5G-Kube, a Kubernetes-native Cloud-native Network Function (CNF) stack that cuts slice instantiation times below 200 ms.
+Larrea et al.
+(2023) developed CoreKube, a cloud-native mobile core whose stateless, message-focused workers are containerized and orchestrated by Kubernetes.
+Additionally, Zhao et al.
+(2025) combined adaptive RAN slicing with hierarchical DRL, achieving E2E KPIs across RAN, TN, and CN.
+Moreover, Wyszkowski and others (2024) present a detailed tutorial that addresses a blind spot in slicing research and standardization: the design-time organization of slice and subnet instances.
 They systematize a taxonomy of slicing entities and formalize the slice/subnet design process, clarifying concepts not uniformly covered in 3GPP or ETSI.
@@
 Compared with other solutions, NASP integrates native B5G support, run-time optimization, and a demonstrable NSaaS prototype, illustrating a viable path toward zero-touch, ZSM-compliant B5G deployments.
+It is worth noting that while the NASP framework architecture is designed to integrate advanced AI-driven closed-loop assurance, we rate its current B5G/AI Ready status as Medium (M) (see Table 1) to reflect that the current prototype relies mainly on rule-based policies, with full AI integration planned for future work.
 4 NSaaS Platform Overview
@@
 The platform mediates between business intent and network implementation by automatically translating high‑level business templates into NS definitions, enabling multi‑domain resource orchestration and monitoring.
-In this context, NASP translates GSMA business templates into versioned Network Slice Templatess through a hierarchical deployment model.
-Business requests collected by the CSMF are transformed into service requirements and forwarded to the NSMF, which assigns slice identifiers (IDs) and delegates domain‑specific tasks.
+In this context, NASP translates GSMA business templates into versioned Network Slice Subnet Templates through a hierarchical deployment model.
+Once initially processed by the Onboard module, business requests are handled by the CSMF, where they are transformed into service requirements and forwarded to the NSMF, which assigns slice identifiers (IDs) and delegates domain‑specific tasks.
 Standardized interfaces (RN1, TN1, and CN1) ensure seamless integration across the RAN, TN, and CN.
@@
 Onboard module
-This NASP module forms the northbound interface through which tenants submit slice requests encoded using the GSMA GST.
+This NASP module acts as the single northbound entry point through which tenants submit their slice requests encoded using the GSMA GST, thereby managing the design-time (or preparation) phase.
+Listing 1 presents a minimal excerpt of a customized eMBB slice request, demonstrating how tenant requirements, such as SLA constraints and domain-specific topological limits, are encoded and submitted.
+Listing 1: Example of a custom tenant slice request encoded as a JSON object.
+1{ "name": "Custom 5G Network Slice",
+2 "NST": {
+3 "type": "custom",
+4 "Slice Attributes": {
+5 "availability": 1,
+6 "Supported Data Network": "internet",
+7 "SSQ": {
+8 "Packet Delay Budget": 0.00012,
+9 "Packet Error Rate": 0.0000001,
+10 "Maximum Data Burts Volume": 0.001},
+11 "UE density": 10000},
+12 "resource_description": {
+13 "core": { "nfs": [{"name": "amf"}, {"name": "smf"}, {"name": "upf"}] },
+14 "ran": { "nfs": [{"name": "ueransim", "type": "gnb", "replicas": 2}] },
+15 "tn": { "routes": [{"name": "backhaul"}] }
+16 }}}
 Incoming service intents first pass through the Requirements micro‑service, where SLA parameters such as latency, throughput, and availability are normalized.
 The data stream is forwarded to Template Matching, whose rule‑based engine correlates the normalized attributes with a 3GPP‑compliant library of versioned NSSTs, selecting the candidate that minimizes resource footprint while maximizing reuse.
-Once a suitable template is chosen, Resource Translations converts logical requirements into concrete descriptors for virtual compute, storage, radio, and transport resources, using mapping rules derived from TMF638; TMF639 and 3gpp.28.541.
+If no existing template mapping can satisfy the normalized attributes, the Onboard module rejects the request and returns an error to the tenant.
+Once the descriptors are prepared and validated, the Onboard module triggers the run-time instantiation phase by forwarding the processed request to the Orchestration module (specifically linking to the CSMF).
+Once a suitable template is chosen, Resource Translations converts logical requirements into concrete descriptors for virtual compute, storage, radio, and transport resources, using mapping rules derived from TM Forum (2024a, b) and 3GPP (2025a).
 The translated artifact is forwarded to the Descriptions service, which renders a standards‑compliant Network Slice Descriptor (NSD) in YAML Ain’t Markup Language (YAML) format, directly consumable by the orchestration tier.
@@
 During onboarding, each GST attribute, such as e2eLatency, serviceAvailability, is persisted as a ServiceSpecification/SlaTarget object via TeleManagement Forum (TMF) 638 (v5.0.0).
-These targets are version‑controlled alongside the NS descriptors and immediately promoted to TMF 640 (v5.0.0) Service Level Objective (SLO) records so that downstream orchestration and assurance functions can reference a single, immutable source of truth.
+These targets are version‑controlled alongside the NS descriptors and immediately promoted to TMF 640 Service Level Objective (SLO) records so that downstream orchestration and assurance functions can reference a single, immutable source of truth.
 Data Management module
@@
 Rather than duplicating data-management functions, the orchestrator treats the Catalog solely as a source of template information and queries the Inventory for the live state of slice resources during provisioning and monitoring.
-The end-to-end workflow, illustrated in Figure 5, begins when the tenant submits a slice request to the CSMF (01); the request is forwarded to the NSMF, which reserves an Single Network Slice Selection Assistance Information (S-NSSAI), retrieves the corresponding slice templates and engages the CN NSSMF to deploy the required CN VNFs under the supervision of the Network Function Virtualization Orchestrator (NFVO) Controller (02–04).
+The end-to-end run-time workflow, illustrated in Figure 5, begins when the tenant’s slice request is forwarded by the Onboard module to the CSMF (01); the request is then forwarded to the NSMF, which reserves an Single Network Slice Selection Assistance Information (S-NSSAI), retrieves the corresponding slice templates and engages the CN NSSMF to deploy the required CN VNFs under the supervision of the Network Function Virtualization Orchestrator (NFVO) Controller (02–04).
 Once an AMF endpoint is available, the RAN NSSMF instantiates the next-generation Node B (gNB) components through the RAN Controller so that they attach to the correct AMF (05–07).
@@
 All configuration artifacts are stored in the Inventory through idempotent transactions, guaranteeing that repeated executions converge to a consistent slice state, and the final slice configuration is preserved for subsequent monitoring and lifecycle management.
-The transformation from high‑level business commitments into enforceable artifacts occurs in four logical phases.
-(i) Production: GST attributes are normalized and stored as TMF 638 (v5.0.0) SlaTarget objects, each of which is linked, via a mapping table inside the Resource Translations micro‑service, to the corresponding TMF 639 (v5.0.0) resource identifiers, e.g., transport bandwidth pools or radio‑spectrum slices.
-(ii) Decomposition: the CSMF converts these targets into TMF 640 (v5.0.0) SLO records and distributes them to the NSMF and the domain‑specific NSSMFs, which embed the derived KPI thresholds in their local controllers.
-(iii) Enforcement: during operation, the Quality‑Assurance loop continuously compares real‑time telemetry with the injected thresholds; any breach raises a TMF 642 (v5.0.0) notification that automatically triggers scaling, relocation, or healing workflows through TMF 645 (v5.0.0) Assurance APIs.
-(iv) Tenant visibility: current SLA status is exposed through a TMF 645 (v5.0.0) /assurance/v1/monitorings endpoint, and tenants may subscribe to NEF webhooks for proactive breach alerts.
+The transformation from high‑level business commitments into enforceable artifacts occurs in four logical phases leveraging TMF Open APIs (v5.0.0).
+(i) Production: GST attributes are normalized and stored as TMF 638 SlaTarget objects, each of which is linked, via a mapping table inside the Resource Translations micro‑service, to the corresponding TMF 639 resource identifiers, e.g., transport bandwidth pools or radio‑spectrum slices.
+(ii) Decomposition: the CSMF converts these targets into TMF 640 SLO records and distributes them to the NSMF and the domain‑specific NSSMFs, which embed the derived KPI thresholds in their local controllers.
+(iii) Enforcement: during operation, the Quality‑Assurance loop continuously compares real‑time telemetry with the injected thresholds; any breach raises a TMF 642 notification that automatically triggers scaling, relocation, or healing workflows through TMF 645 Assurance APIs.
+(iv) Tenant visibility: current SLA status is exposed through a TMF 645 /assurance/v1/monitorings endpoint, and tenants may subscribe to NEF webhooks for proactive breach alerts.
 The orchestration workflow follows the 3GPP management and service specifications, including TS 28.531 and TS 28.541 for slice lifecycle coordination, as well as TS 23.501 and TS 38.413 for inter-domain reference points.
@@
 This module supervises every active NS through a non‑real‑time control loop that begins with Telemetry Discovery, which retrieves multilayer observability data (counters, alarms, traces, and logs) via the NEF, RAN monitoring interfaces, and SDN probes.
-Although transport-domain SDN telemetry (OpenFlow and sFlow Protocol (sFlow)/ IP Flow Information Export (IPFIX)) is not natively covered by 3GPP TS 28 5xx, a lightweight adaptor inside the Data Translation micro-service enriches every record with the S-NSSAI and converts raw counters, such as port-tx_bytes, queue delay, packet-drop events, into their 3GPP-compliant counterparts, defined in 3gpp.28.552.
+Although transport-domain SDN telemetry (OpenFlow and sFlow Protocol (sFlow)/ IP Flow Information Export (IPFIX)) is not natively covered by 3GPP TS 28 5xx, a lightweight adaptor inside the Data Translation micro-service enriches every record with the S-NSSAI and converts raw counters, such as port-tx_bytes, queue delay, packet-drop events, into their 3GPP-compliant counterparts, defined in 3GPP (2025d).
 The normalized KPIs are pushed (i) towards the NWDAF for analytics or (ii) as TMF 642 events for fault management.
@@
 For example, RN1 submits a RAN Slice Set to the Kubernetes (K8s) API server managing the cloud‑native RAN.
-The manifest is deployed through Helm releases, Network Attachment Definitions, and accompanying custom resources configuring the required gNB functions.
-Moreover, TN1 delivers a Path Intent over a RESTful Hypertext Transfer Protocol (HTTP)/2 channel to the SDN controller, translating the request into OpenFlow 1.5 flow‑mod messages.
+The manifest is deployed via Helm releases, Network Attachment Definitions, and accompanying custom resources that configure the required gNB functions.
+Moreover, TN1 delivers a Path Intent over a RESTful Hypertext Transfer Protocol (HTTP)/2 channel to the SDN controller.
+Because this intent is technology-agnostic, the controller can translate it into the appropriate southbound protocols based on the underlying network equipment.
+For instance, it may employ OpenFlow 1.5 flow‑mod messages for fine-grained traffic engineering on virtual switches, or push YANG-modeled configurations via Netconf to manage physical transport routers.
 Finally, CN1 forwards a CN Slice Set to the K8s API server that orchestrates the CN.
@@
 Each pod’s network interfaces are dynamically reconfigured to route traffic using VLAN-based segmentation, providing isolation and meeting performance requirements (Figure 7).
-Distinct VLAN IDs enforce specific QoS policies: a short path prioritizes low latency, while a long path provides redundancy.
-Although the TN is programmed via ONOS/OpenFlow, we keep a lightweight 802.1Q tag on every packet.
+Although the TN is programmed via ONOS/OpenFlow, every packet retains a lightweight 802.1Q tag.
 Distinct VLAN IDs encode slice-specific QoS policies: a short path prioritizes low latency for URLLC, whereas a longer path provides redundancy for mMTC and Shared slices.
-Although the TN is programmed via ONOS/OpenFlow, packets retain lightweight 802.1Q tags.
 These tags act as in-band slice and QoS markers, allowing switches to differentiate flows locally, reduce flow-table pressure, and perform deterministic path selection before any controller interaction.
-This mechanism preserves layer-2 isolation and remains compatible with legacy Ethernet segments (IEEE8021Q-2018).
+This mechanism preserves layer-2 isolation and remains compatible with legacy Ethernet segments (IEEE, 2018).
 Figure 7: Emulated TN topology with VLAN-based segmentation.
@@
 Quality Assurance Closed Loop module: Implements telemetry collection (Prometheus exporters), KPI extraction (deployment time, attach latency, availability, resource consumption), and threshold- or rule-based corrective actions such as AMF reconfiguration.
-Advanced AI-driven analytics (e.g., predictive scaling, anomaly classification) are currently placeholders for future work.
+Although the architecture includes components for AI-driven analytics, the current implementation relies primarily on rule-based mechanisms, leaving advanced AI capabilities (e.g., predictive scaling, anomaly classification) for future work.
 Interfaces: Implemented (RN1, TN1, CN1) as a unified northbound REST layer that normalizes slice intents into domain‑specific manifests (Helm/YAML for Kubernetes, JSON payloads for ONOS).
@@
 Service Scenarios and Network-Function Mapping
-We instantiated four slice scenarios to produce the results discussed later: (i) URLLC, (ii) mMTC (a.k.a., mIoT in later figures), (iii) Shared (an eMBB-type slice that reuses existing control‑plane Network Functions to measure best‑case provisioning), and (iv) non‑3GPP access (offload via N3IWF).
+We instantiated four slice scenarios to produce the results discussed later: (i) URLLC, (ii) mMTC (a.k.a., mIoT in later figures), (iii) Shared (an eMBB-type slice that reuses existing control‑plane NFs to measure best‑case provisioning), and (iv) non‑3GPP access (offload via N3IWF).
 Table 2 summarizes the architectural deltas that drive the different deployment and performance outcomes.
@@
 Testbed Infrastructure and Network Topology
-Experiments were executed on Virtual Machines provided by DigitalOcean.
-The testbed comprises heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0-88-generic.
-Control-plane components were deployed on s-4vcpu-8gb instances (4 vCPU, 8 GB RAM), whereas data-plane components used s-8vcpu-16gb instances (8 vCPU, 16 GB RAM).
-Dedicated VMs were allocated to Open Network Automation Platform (ONAP), Mininet with ONOS, and the Kubernetes cluster.
-The Kubernetes deployment (v1.28.3) used Calico (Calico) and Multus (Multus) to support multi-tenancy, with autoscaling thresholds configured at 80% for scale-up and 20% for scale-down.
-Persistent monitoring data were exported to a Prometheus–Grafana stack.
-Experiments were conducted on VMs provided by DigitalOcean.
-The cloud testbed consists of heterogeneous VMs running Ubuntu 22.04 LTS with Linux kernel 5.15.0‑88‑generic.
+Experiments were conducted on Virtual Machines provided by DigitalOcean.
+The cloud testbed comprises heterogeneous VMs running Ubuntu 22.04 LTS with the Linux kernel 5.15.0‑88‑generic.
 Control‑plane components ran on s‑4vcpu‑8gb instances (4 Virtual CPU (vCPU), 8 GB RAM), whereas data‑plane elements used s‑8vcpu‑16gb instances (8 vCPU, 16 GB RAM).
-We dedicated a separate VMs to ONAP, Mininet with ONOS, and a Kubernetes cluster.
-We used Kubernetes v1.28.3 with Calico and Multus to support multi‑tenancy, with autoscaling thresholds of 80% (scale‑up) and 20% (scale‑down).
+We dedicated a separate VMs to Open Network Automation Platform (ONAP), Mininet with ONOS, and a Kubernetes cluster.
+We used Kubernetes v1.28.3 with Calico (Calico) and Multus (Multus) to support multi‑tenancy, with autoscaling thresholds of 80% (scale‑up) and 20% (scale‑down).
 Persistent monitoring data were exported to a Prometheus–Grafana stack, and all timestamps were converted to Coordinated Universal Time (UTC) to avoid clock‑skew artifacts.
-In this work, we propose three types of cloud sites for the network performance evaluation: Regional (Central), Metropolitan (Edge), and Internal (Extreme‑Low‑Latency Edge), as defined by gustavo2024evaluating.
+In this work, we propose three types of cloud sites for the network performance evaluation: Regional (Central), Metropolitan (Edge), and Internal (Extreme‑Low‑Latency Edge), as defined by Bruno and others (2024).
 The Internal site offered limited instance types, so we normalized cost comparisons using pricing from all sites.
@@
 The results are organized along two dimensions: (i) the design and deployment of E2E NSIs and (ii) their performance and lifecycle management.
+In addition to the within-testbed measurements reported in this section, we include a structured cross-study quantitative contextualization.
+This analysis connects the Related Work discussion with the NASP experimental results.
+The cited solutions differ substantially in architectural scope (e.g., CN-focused versus E2E RAN/TN/CN orchestration), lifecycle coverage, and experimental setup (e.g., simulation, emulation, real, or hybrid testbeds).
+Therefore, we discuss representative reported metrics from prior work and preserve their original definitions and units.
+Furthermore, we qualify each comparison according to its degree of alignment with NASP (Direct, Partial, Contextual, or Not comparable).
+This approach provides quantitative grounding without forcing equivalence when metric semantics, control boundaries, or measurement scope differ.
 Figure 9: Slice deployment time across four scenarios (URLLC, mMTC, Shared eMBB reuse, non-3GPP offload).
+7.1 Slice Provisioning and Instantiation Delay
 The NASP design-and-deployment evaluation quantifies the time required for slice provisioning.
@@
 These observations reinforce that streamlined control domains, such as TN, benefit from specialized controllers that minimize latency and improve overall efficiency of deployment.
+To further contextualize the provisioning results, we relate NASP’s E2E slice deployment time (22–53 s, Fig.
+9) and stepwise orchestration breakdown (Fig.
+10) to timings from prior studies with different scopes.
+The sub-200 ms instantiation reported by Scotece and others (2023) refers to localized cloud-native CNF/core startup.
+This evaluation serves only as a contextual lower-bound reference, not a direct E2E comparator.
+In contrast, Dalgitsis and others (2024) reports strategy-dependent slice deployment times of 27.0, 19.4, 21.9, $\approx 4$, and $\approx 4$ s in a federated multi-AD cloud-native 5G setup (post-federation deployment under different NF sharing/isolation strategies), which provides a partial comparator for NASP’s 22–53 s E2E provisioning.
+Quantitatively, these results place both studies in the seconds-to-tens-of-seconds range when orchestration scope extends beyond isolated function startup, and show that deployment time varies materially with workflow structure and sharing/isolation choices.
+We use the 33–38 ms (London) and 45 ms (Ireland) cloud-region latency values reported by Dalgitsis and others (2024) only as environmental context for edge-cloud experiments, not as direct slice-KPI outcomes, to avoid conflating network-path conditions with NASP’s service-level latency measurements.
+7.2 Lifecycle Management and Dynamic Reconfiguration
 The slice-lifecycle evaluation investigates UE attachment latency, dynamic slice reconfiguration, and resource utilization.
@@
 No successful registration responses are observed between 23 s and 47 s, the exact interval highlighted by the “Timeout” arrow, because all attempts expire while the control plane is rebuilding.
+We contextualize the runtime behavior observed in NASP against representative cloud-native slicing and core-network studies, while preserving the original metric semantics of those works.
+Zhao et al.
+(2025) reports $64.2\%$ total operating-cost reduction and $45.5\%$ normalized-performance improvement, as well as converged orchestration outcomes of $4.3/1.79$ (cost / normalized performance) versus $12/1.23$ for Atlas, and dynamic adaptation after slice departure with $25.6\%$ cost reduction and $12.8\%$ performance improvement in 1 slot.
+These results provide a partial/contextual reference for NASP’s runtime adaptation, cost-efficiency, and service-quality-oriented evaluation dimensions.
+Zhao et al.
+(2025) also reports up to $53.9\%$ degradation (normalized performance) for comparison systems beyond 3 slices (up to 5 slices), which we use as a partial scalability reference for multi-slice concurrency effects, although the slice abstractions and resource models differ from NASP Zhao et al.
+(2025).
+Likewise, CoreKube (Larrea et al.
+(2023)) reports control-plane scalability thresholds where NextEPC saturates beyond $\approx 12$ requests/s, Open5GS saturates after $\approx 120$ requests/s, and CoreKube scales beyond; a burst-handling result of 560 msgs/s from 250 UEs; resilience experiments with a 200 s critical-error interval (about $\sim 20\times$ more frequent than the referenced baseline); and AMF CPU imbalance of $91.4/1.4/1.4\%$ across three AMF pods in the K8s-Open5GS case.
+These results provide contextual evidence for CN/control-plane scalability and robustness, and quantitatively demonstrate that replication alone does not ensure balanced scaling without adequate traffic distribution and orchestration.
+Finally, we explicitly note that the disruption interval discussed in the comparison setup corresponds to the baseline environment without NASP and is not attributed to NASP-managed AMF redeployment.
+This part of the discussion presents a contextual comparison of reconfiguration/failure models and automation behavior rather than a direct outage benchmark.
 Figure 13: Registration latency during the reconfiguration process.
@@
 Figure 14: Slice utilization before and after dynamic reconfiguration (admission control event at dashed line).
+7.3 Resource Utilization and Cost Analysis
 We also monitored resource utilization continuously during slice deployment.
@@
 Monthly prices were normalized by advertised resources so that the regression coefficients express the marginal cost, in $\mathdollar\,\text{month}^{-1}$, of one vCPU and one GiB of RAM.
-Table 3: Description of Edge, Metropolitan, and Central tier instances (amazonStartedWith).
+Table 3: Description of Edge, Metropolitan, and Central tier instances (Amazon, 2024).
 Type Size vCPU RAM (GB) Storage (GB) Price/Month
@@
 Figure 16: Predicted monthly cost as a function of vCPU and RAM utilization.
-Overall, these experimental results validate the capability of the NASP architecture to efficiently support network‑slice design, deployment, and dynamic lifecycle management.
+These experimental results validate the capability of the NASP architecture to efficiently support network‑slice design, deployment, and dynamic lifecycle management.
 Beyond demonstrating flexibility across four 5G scenarios, the evaluation indicates a balanced trade‑off between deployment speed, run‑time performance, and operational cost.
-In this context, NASP provide a viable platform for next‑generation network slicing.
+Note that our current evaluation environment uses the my5G-RANTester for RAN emulation, with a focus on end-to-end orchestration workflows.
+This is because NASP operates at the orchestration layer, managing the lifecycle at the network-function level rather than at the individual-flow or physical-link level.
+Mobility and handover procedures are executed within the RAN execution layer, whereas NASP interfaces with RAN components at the management and orchestration layer.
+While mobility and large-scale attachment effects influence runtime service performance and QoS metrics, they do not affect the architectural validity of the orchestration mechanisms evaluated in this work.
+In this context, NASP provides a viable platform for next‑generation network slicing.
+The experiments consider up to 250 concurrent UEs per slice, and our experimental topology cannot fully capture topology-driven effects such as distributed domain coordination, transport-domain expansion, and end-to-end operational delays.
+This is because the evaluation focuses on the scalability and efficiency of the control plane and the orchestration lifecycle (e.g., slice instantiation, configuration, and end-to-end lifecycle management) in a controlled proof-of-concept environment, rather than on stressing data-plane capacity typical of large-scale mMTC deployments.
+In our framework, the orchestration overhead is primarily determined by the number of managed slices, instantiated network functions, and lifecycle events, rather than directly by the number of attached devices or the raw number of switches or physical links, although the scale and heterogeneity of the topology may still influence the practical orchestration latency.
+Therefore, the selected experimental scale enables us to capture the relevant behavior of the orchestration mechanisms as slice management complexity increases.
+While mMTC scenarios typically involve much larger device populations, the goal of this study is to evaluate the orchestration framework’s ability to efficiently deploy, manage, and coordinate multiple network slices.
+Future work will extend the evaluation to larger device populations and heterogeneous, geographically distributed topologies, including real-world RAN environments and mobility scenarios, to further assess the system under large-scale mMTC conditions.
 8 Conclusion
@@
 Beyond the architectural contributions, the seamless integration of standardized GSMA templates, the unification of definitions across major mobile network entities to realize an E2E architecture, and the capability to instantiate network slices for Non-3GPP (non3GPP) applications open several avenues for future work.
-Planned efforts include: (a) integrating a physical RAN testbed to eliminate access-layer abstractions, (b) scaling the orchestration logic to support many concurrent slices, (c) embedding security and privacy assurance loops, (d) validating interoperability with commercial MANO stacks, such as ONAP and OSM, while aligning interfaces with emerging Open Radio Access Network (O-RAN) specifications in multi-operator environments, and (e) enhancing mobility management for seamless session continuity, particularly in light of the observed slice-establishment times.
+Planned efforts include: (a) integrating real-world physical RAN deployments to assess radio-level effects and mobility procedures that were abstracted in this study, (b) scaling the physical topology across heterogeneous domains to stress-test the data plane under massive machine-type communications (mMTC) limitations, (c) embedding advanced AI-driven analytics for predictive scaling and anomaly classification within the closed-loop assurance mechanisms, and (d) validating interoperability with commercial MANO stacks, such as ONAP and OSM, while aligning interfaces with emerging Open Radio Access Network (O-RAN) specifications in multi-operator environments.
 Acknowledgment